### v4

# Trilinear characterizations of the Fourier extension conjecture on the paraboloid in three dimensions

Abstract.
We first prove that a local trilinear extension inequality on the paraboloid in $\mathbb{R}^{3}$ is equivalent to the Fourier restriction conjecture in $\mathbb{R}^{3}$, which may be viewed as a companion result to either the transversal trilinear inequality of Bennett, Carbery and Tao in [BeCaTa], or to the bilinear characterization due to Tao, Vargas and Vega in [TaVaVe]. Namely, we prove that the paraboloid restriction conjecture holds in $\mathbb{R}^{3}$ if and only if for every $q>3$ there is $\nu>0$ such that the trilinear inequality,
$\left\|\mathcal{E}f_{1}\mathcal{E}f_{2}\mathcal{E}f_{3}\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}\leq C_{\nu}R^{\varepsilon}\mathop{\displaystyle\prod}\limits_{k=1}^{3}\left\|f_{k}\right\|_{L^{\infty}\left(U\right)}\ ,\ \ \ \ \ \text{for }R>1,\varepsilon>0$
holds when taken over all $f_{k}\in L^{\infty}\left(U_{k}\right)$, and all triples $\left(U_{1},U_{2},U_{3}\right)$ of squares that satisfy the $\nu$-disjoint condition
$\mathop{\rm diam}\left[\Phi\left(U_{k}\right)\right]\approx\nu\text{ and }\mathop{\rm dist}\left[\Phi\left(U_{k}\right),\Phi\left(U_{j}\right)\right]\geq\nu,\text{ for }1\leq j,k\leq 3,$
where $\Phi$ is the standard parameterization of the paraboloid $\mathbb{P}^{2}$ and $\mathcal{E}f=\left[\Phi_{\ast}\left(f\left(x\right)dx\right)\right]^{\wedge}$ is the associated Fourier extension operator. The proof follows an argument of Bourgain and Guth [BoGu], but exploiting the fact that the problematic Case 3 in their argument dissolves due to the $\nu$-disjoint assumption, as opposed to the $\nu$-transversal assumption in [BeCaTa].
Then we prove the equivalence of the above trilinear Fourier extension conjecture with the special case of testing a local trilinear inequality over certain smooth Alpert pseudoprojections $\mathsf{Q}_{s,U}^{\eta}f$, representing the weakest such inequality equivalent to the Fourier extension conjecture that the authors could find. This special inequality has the following form. For every $q>3$ there is $\nu>0$ such that
$\left(\int_{B\left(0,2^{s}\right)\setminus B\left(0,2^{s-1}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \lesssim 2^{\varepsilon s}\left\|f_{1}\right\|_{L^{\infty}}\left\|f_{2}\right\|_{L^{\infty}}\left\|f_{3}\right\|_{L^{\infty}}\ ,$
for all $\nu$-disjoint triples $\left(U_{1},U_{2},U_{3}\right)$, all $\varepsilon>0$, $s\in\mathbb{N}$ and $s_{1}\leq s_{2}\leq s_{3}$ with $s_{2},s_{3}$ arbitrarily close to $s$.
The extension to other quadratic surfaces of positive Gaussian curvature in $\mathbb{R}^{3}$ is straightforward.
Eric Sawyer’s research supported in part by a grant from the National Sciences and Engineering Research Council of Canada
1. Introduction
In this paper we show that a variant of the trilinear Fourier extension inequality of Bennett, Carbery and Tao [BeCaTa] [^1] See [Tao2] for a sharpening of these results. is equivalent to the Fourier extension conjecture in dimension three, at least for quadratic surfaces of positive Gaussian curvature, which include the paraboloid. This can also be viewed as a companion result to the bilinear characterizations in Tao, Vargas and Vega [TaVaVe]. Of course the trilinear characterization implies the corresponding bilinear one by Hölder’s inequality222applied with exponents $\left(3,3,3\right)$ to the factorization $\left|\mathcal{E}f_{1}\mathcal{E}f_{2}\mathcal{E}f_{3}\right|=\sqrt{\left|\mathcal{E}f_{1}\mathcal{E}f_{2}\right|}\sqrt{\left|\mathcal{E}f_{2}\mathcal{E}f_{3}\right|}\sqrt{\left|\mathcal{E}f_{3}\mathcal{E}f_{1}\right|}.$ It is possible that the bilinear proof in [TaVaVe] can be adapted to the disjoint trilinear setting, but the Whitney condition in three dimensions complicates matters, and we will not pursue this here..
We will weaken the transversality hypothesis in [BeCaTa] to a disjoint hypothesis (while retaining the embeddings into the paraboloid), as well as weakening the bound in the conclusion, and then adapt an argument of Bourgain and Guth [BoGu] to derive the Fourier extension conjecture as a consequence of this disjoint Fourier trilinear inequality. The key point here is that the problematic Case 3 in the argument of [BoGu, Section 2] is eliminated, along with their restriction to $p>\frac{10}{3}$, by the disjoint assumption.
In Theorem 10 in the third section of the paper, we state and prove another characterization of the Fourier extension inequality, in which we restrict the functions in the disjoint trilinear conjecture to special smooth Alpert pseudoprojections, representing the ‘simplest’ characterization that the authors could find. In the appendix immediately following that, we sketch how a square function modification of the arguments used here could give an alternate, and arguably simpler, proof of the probabilistic Fourier extension theorem in [Saw7] for the paraboloid in three dimensions [^3] The result in [Saw7] was stated only for the sphere, but easily extends to smooth compact surfaces with positive Hessian..
We will only give details of proofs for our results in the special case of the paraboloid $\mathbb{P}^{2}$ in three dimensions here, since the arguments are virtually the same for quadratic surfaces with positive Hessian in $\mathbb{R}^{3}$. Denote the Fourier extension operator $\mathcal{E}$ by,
$\mathcal{E}f\left(\xi\right)\equiv\left[\Phi_{\ast}\left(f\left(x\right)dx\right)\right]^{\wedge}\left(\xi\right),\ \ \ \ \ \text{for }\xi\in\mathbb{R}^{3},$
where $\Phi_{\ast}\left(f\left(x\right)dx\right)$ denotes the pushforward of the measure $f\left(x\right)dx$ supported in $U\subset B_{\mathbb{R}^{2}}\left(0,\frac{1}{2}\right)$ to the paraboloid $\mathbb{P}^{2}$ under the usual parameterization $\Phi:U\rightarrow\mathbb{P}^{2}$ by $\Phi\left(x\right)=\left(x_{1},x_{2},x_{1}^{2}+x_{2}^{2}\right)$ for $x=\left(x_{1},x_{2}\right)\in U$.
Conjecture 1 (Fourier extension).
The Fourier extension conjecture for the paraboloid $\mathbb{P}^{2}$ in $\mathbb{R}^{3}$ is the assertion that
(1.1) $\left\|\mathcal{E}f\right\|_{L^{q}\left(\mathbb{R}^{3}\right)}\lesssim\left\|f\right\|_{L^{q}\left(U\right)},\ \ \ \ \ \text{for }q>3.$
which we also denote by $\mathcal{E}\left(\otimes_{1}L^{q}\rightarrow L^{q}\right)$.
1.1. Statements of main theorems
The analogous Fourier extension conjecture for the $n$-dimensional sphere $\mathbb{S}^{n-1}$ was made in 1967 by E. M. Stein, see e.g. [Ste2, see the Notes at the end of Chapter IX, p. 432, where Stein proved the restriction conjecture for $1\leq p<\frac{4n}{3n+1}$] and [Ste]. The two-dimensional case of the Fourier extension conjecture was proved over half a century ago by L. Carleson and P. Sjölin [CaSj], see also C. Fefferman [Fef] and A. Zygmund [Zyg]. A web search reveals much progress on extension theorems, as well as Kakeya theorems, in the ensuing years. In particular, the Kakeya set conjecture has recently been proved in dimension $n=3$ by Hong Wang and Joshua Zahl [WaZa].
Here is our weakening of the transversality condition as introduced by Muscalu and Oliveira [MuOl], which we refer to as a disjoint condition.
Definition 2.
Let $\varepsilon,\nu>0$, $1<q<\infty$ and $q\leq p\leq\infty$. Denote by $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{p}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ the trilinear Fourier extension inequality
(1.2) $\left\|\mathcal{E}f_{1}\ \mathcal{E}f_{2}\ \mathcal{E}f_{3}\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}\leq\left(C_{\varepsilon,\nu,p,q}R^{\varepsilon}\right)^{3}\left\|f_{1}\right\|_{L^{p}\left(U_{1}\right)}\left\|f_{2}\right\|_{L^{p}\left(U_{2}\right)}\left\|f_{3}\right\|_{L^{p}\left(U_{3}\right)}\ ,$
taken over all $R\geq 1$, $f_{k}\in L^{p}\left(U_{k}\right)$, and all triples $\left(U_{1},U_{2},U_{3}\right)\subset U^{3}$ that satisfy the $\nu$-disjoint condition,
(1.3) $\mathop{\rm diam}\left[\Phi\left(U_{k}\right)\right]\approx\nu\text{ and }\mathop{\rm dist}\left[\Phi\left(U_{k}\right),\Phi\left(U_{j}\right)\right]\geq\nu,\text{ for }1\leq j,k\leq 3,$
and where the constant $C_{\varepsilon,\nu,p,q}$ is independent of $R\geq 1$ and the functions $f_{k}\in L^{p}\left(U_{k}\right)$.
Note that (1.2) is invariant under translation of the ball $B\left(0,R\right)$ in $\mathbb{R}^{3}$.
Our first theorem is the equivalence of Fourier extension and disjoint trilinear Fourier extension.
Theorem 3.
The Fourier extension conjecture holds for the paraboloid $\mathbb{P}^{2}$ in $\mathbb{R}^{3}$ if and only if for every $q>3$ there is $\nu>0$ such that the disjoint trilinear inequality $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{q}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ holds for all $\varepsilon>0$. More generally, the following statements are equivalent:
  1. (1)
    $\mathcal{E}\left(\otimes_{1}L^{q}\rightarrow L^{q}\right)$ for all $q>3$,
  2. (2)
    For every $q>3$ there is $\nu>0$ such that $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{q}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ for all $\varepsilon>0$,
  3. (3)
    For every $q>3$ there is $\nu>0$ such that $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ for all $\varepsilon>0$.
The proof shows that we can take $\nu=O\left(2^{-\frac{3q}{q-3}}\right)$, see (2.22). We exploit the fact that the argument in [BoGu, See the Remark after (6.19) on page 1248.] simply requires a trilinear inequality for a sufficiently small separation constant $\nu>0$, depending only on $q>3$. This theorem also gives a characterization of the Bôchner-Riesz inequality on the paraboloid (and also on quadratic surfaces of positive Gaussian curvature as mentioned above) in view of the equivalence of Fourier extension and Bôchner-Riesz for such surfaces, see Carbery [Car] and [Tao] for this.
Remark 4.
It is not clear that the Fourier extension inequality for the sphere can be characterized by the methods used here, since when considering (1.1) for a surface $S$ without invariance under a quadratic dilation, such as the sphere, the proof of Case 2 below requires introduction of perturbations $S^{\prime}$ of the surface $S$ as well, which then prevents a characterization of (1.1) in terms of (1.2) - see [TaVaVe, page 297] where this is discussed. Nevertheless, we conjecture that the methods used here can be adapted to prove that (1.1) holds for all surfaces $S$ in $\mathbb{R}^{3}$ of positive Gaussian curvature that are bounded by some fixed constant $A$, if and only if (1.2) holds uniformly for the same class of surfaces444the reason being that the parabolic rescalings used in the proof do not exit the class of surfaces under consideration..
Theorem 3 can be put into context as follows. In the special case where the patches $\Phi\left(U_{1}\right),\Phi\left(U_{2}\right),\Phi\left(U_{3}\right)$ are $\nu$-transverse, then the trilinear inquality (1.2) is the trilinear inequality proved by Bennett, Carbery and Tao [BeCaTa]. In the more general case when the patches $\Phi\left(U_{1}\right),\Phi\left(U_{2}\right),\Phi\left(U_{3}\right)$ are merely assumed $\nu$-disjoint, then the trilinear inequality (1.2) implies the Fourier extension conjecture.
1.2. Easy directions of the proof
The implication $\left(1\right)\Longrightarrow\left(2\right)$ of Theorem 3 follows from applying Hölder’s inequality with exponents $\frac{1}{3},\frac{1}{3},\frac{1}{3}$ to show that (1.1) implies (1.2) with $\varepsilon=0$ and even without the $\nu$-disjoint condition (1.3):
$\left(\int_{\mathbb{R}^{3}}\left|\mathcal{E}f_{1}\left(\xi\right)\ \mathcal{E}f_{2}\left(\xi\right)\ \mathcal{E}f_{3}\left(\xi\right)\right|^{\frac{q}{3}}d\xi\right)^{\frac{3}{q}}\lesssim\left(\int_{\mathbb{R}^{3}}\left|\mathcal{E}f_{1}\left(\xi\right)\right|^{q}d\xi\right)^{\frac{1}{q}}\left(\int_{\mathbb{R}^{3}}\left|\mathcal{E}f_{2}\left(\xi\right)\right|^{q}d\xi\right)^{\frac{1}{q}}\left(\int_{\mathbb{R}^{3}}\left|\mathcal{E}f_{3}\left(\xi\right)\right|^{q}d\xi\right)^{\frac{1}{q}}$
$\lesssim\left(\int_{U}\left|f_{1}\left(x\right)\right|^{q}dx\right)^{\frac{1}{q}}\left(\int_{U}\left|f_{2}\left(x\right)\right|^{q}dx\right)^{\frac{1}{q}}\left(\int_{U}\left|f_{3}\left(x\right)\right|^{q}dx\right)^{\frac{1}{q}}=\left\|f_{1}\right\|_{L^{q}}\left\|f_{2}\right\|_{L^{q}}\left\|f_{3}\right\|_{L^{q}}\ .$
The implication $\left(2\right)\Longrightarrow\left(3\right)$ follows from the embedding $\left\|f_{k}\right\|_{L^{q}\left(U_{k}\right)}\leq\left\|f_{k}\right\|_{L^{\infty}\left(U_{k}\right)}\left|U_{k}\right|^{\frac{1}{q}}\leq\left\|f_{k}\right\|_{L^{\infty}\left(U_{k}\right)}$.
The next section of this paper is devoted to an adaptation of the argument of Bourgain and Guth [BoGu, Section 2] that will show that $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ for all $q>3$ and $\varepsilon>0$, implies the Fourier extension conjecture $\mathcal{E}\left(\otimes_{1}L^{q}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ for $q>3$, i.e. (1.1) in $\mathbb{R}^{3}$, thereby establishing the implication $\left(3\right)\Longrightarrow\left(1\right)$, and completing the proof of Theorem 3. In the third section we obtain the equivalence of the Fourier extension conjecture with an Alpert disjoint trilinear conjecture, and in the final appendix section, we sketch an alternate proof of the probabilistic Fourier extension theorem in [Saw7].
2. Proof that disjoint trilinear extension implies Fourier extension
A natural approach to proving $\left(3\right)\Longrightarrow\left(1\right)$ is to write $f=\sum_{K\in\mathcal{G}_{\lambda}\left[U\right]}\mathbf{1}_{K}f$, where $\nu=2^{-\lambda}$ and the squares $K\in\mathcal{G}_{\lambda}\left[U\right]$ tile $U$ and have side length $2^{-\lambda}$. Then we have
$\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,R\right)\right)}^{q}$ $=\left\|\left(\sum_{K\in\mathcal{G}_{\lambda}\left[U\right]}\mathcal{E}\left(\mathbf{1}_{K}f\right)\right)^{3}\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}^{\frac{q}{3}}=\left\|\sum_{\left(K_{1},K_{2},K_{3}\right)\in\mathcal{G}_{\lambda}\left[U\right]^{3}}\mathcal{E}\left(\mathbf{1}_{K_{1}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{2}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{3}}f\right)\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}^{\frac{q}{3}}$
$=\left\|\left\{\sum_{\left(K_{1},K_{2},K_{3}\right)\in\Gamma_{1}}+\sum_{\left(K_{1},K_{2},K_{3}\right)\in\Gamma_{2}}+\sum_{\left(K_{1},K_{2},K_{3}\right)\in\Gamma_{3}}\right\}\mathcal{E}\left(\mathbf{1}_{K_{1}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{2}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{3}}f\right)\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}^{\frac{q}{3}}$
$\lesssim\sum_{\alpha=1}^{3}\left\|\sum_{\left(K_{1},K_{2},K_{3}\right)\in\Gamma_{\alpha}}\mathcal{E}\left(\mathbf{1}_{K_{1}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{2}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{3}}f\right)\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}^{\frac{q}{3}}\equiv\sum_{\alpha=1}^{3}T_{\alpha}\ ,$
where
$\Gamma_{1}$ $\equiv\left\{\left(K_{1},K_{2},K_{3}\right)\in\mathcal{G}_{\lambda}\left[U\right]^{3}:\text{no pair of squares touch}\right\},$
$\Gamma_{2}$ $\equiv\left\{\left(K_{1},K_{2},K_{3}\right)\in\mathcal{G}_{\lambda}\left[U\right]^{3}:\text{for exactly one pair of squares touch}\right\},$
$\Gamma_{3}$ $\equiv\left\{\left(K_{1},K_{2},K_{3}\right)\in\mathcal{G}_{\lambda}\left[U\right]^{3}:\text{every square touches another}\right\}.$
Term $T_{1}$ can be controlled by $C_{\varepsilon,\nu,\infty,q}R^{\varepsilon}$ using (1.2). Term $T_{3}$ can be controlled using parabolic rescaling as in Case 2 below,
$\left\|\sum_{\left(K_{1},K_{2},K_{3}\right)\in\Gamma_{3}}\mathcal{E}\left(\mathbf{1}_{K_{1}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{2}}f\right)\ \mathcal{E}\left(\mathbf{1}_{K_{3}}f\right)\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}^{\frac{q}{3}}\lesssim\left\|\sum_{K\in\mathcal{G}_{\lambda}\left[U\right]}\mathcal{E}\left(\mathbf{1}_{K}f\right)^{3}\right\|_{L^{\frac{q}{3}}\left(B\left(0,R\right)\right)}^{\frac{q}{3}}$
$\leq\left(\#\mathcal{G}_{\lambda}\left[U\right]\right)^{\frac{q}{3}-1}\sum_{K\in\mathcal{G}_{\lambda}\left[U\right]}\int_{B\left(0,R\right)}\left|\mathcal{E}\left(\mathbf{1}_{K}f\right)\left(\xi\right)\right|^{q}d\xi\lesssim\left(\#\mathcal{G}_{\lambda}\left[U\right]\right)^{\frac{q}{3}-1}\sum_{K\in\mathcal{G}_{\lambda}\left[U\right]}\left(2^{-\lambda}\right)^{\left(2q-4\right)}\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,2^{-\lambda}R\right)\right)}^{q}$
$\leq\left(\#\mathcal{G}_{\lambda}\left[U\right]\right)^{\frac{q}{3}}\left(2^{-\lambda}\right)^{2q-4}\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,2^{-\lambda}R\right)\right)}^{q}=\left(2^{-\lambda}\right)^{\frac{4}{3}q-4}\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,2^{-\lambda}R\right)\right)}^{q},$
which is at most $\frac{1}{2}\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,R\right)\right)}^{q}$ if $\lambda$ is chosen sufficiently large depending on $q>3$. Then we have
$\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,R\right)\right)}^{q}$ $\leq T_{1}+T_{2}+T_{3}\leq C_{\varepsilon,\nu,\infty,q}R^{\varepsilon}+T_{2}+\frac{1}{2}\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,R\right)\right)}^{q},$
$\Longrightarrow\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,R\right)\right)}^{q}\leq C_{\varepsilon,\nu,\infty,q}R^{\varepsilon}+T_{2}\ ,$
but unfortunately, term $T_{2}$ is problematic since the same argument produces a larger power $\left(\#\mathcal{G}_{\lambda}\left[U\right]\right)^{\frac{q}{3}+1}$ due to summing over two independent squares in $\mathcal{G}_{\lambda}\left[U\right]$. The resulting estimate $\left(2^{-\lambda}\right)^{\left(2q-8\right)}\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,R\right)\right)}^{q}$ cannot be absorbed unless $q>4$.
Here we will use the disjoint trilinear estimate $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ to essentially eliminate the difficult Case 3 of the Bourgain and Guth argument in [BoGu, Section 2], along with the restriction $p>\frac{10}{3}$ there. This results in an optimal local linear inequality, which in turn proves the Fourier extension conjecture in three dimensions by Nikishin-Maurey-Pisier theory and $\varepsilon$-removal techniques.
Suppose $S$ is a compact smooth hypersurface contained in $\mathbb{R}^{3}$ that is contained in the paraboloid $\mathbb{P}^{2}$, and denote surface measure on $S$ by $\sigma$. The next definition is specialized from [BoGu].
Definition 5.
For $1<q<\infty$ and $R>0$ define $Q_{R}^{\left(q\right)}$ to be the best constant in the local linear Fourier extension inequality,
$\left(\int_{B\left(0,R\right)}\left|\widehat{\Phi_{\ast}f}\left(\xi\right)\right|^{q}d\xi\right)^{\frac{1}{q}}\leq Q_{R}^{\left(q\right)}\left\|f\right\|_{L^{\infty}\left(U\right)}\ ,$
i.e.
(2.1) $Q_{R}^{\left(q\right)}\equiv\sup_{\left\|f\right\|_{L^{\infty}\left(U\right)}\leq 1}\left(\int_{B\left(0,R\right)}\left|\mathcal{E}f\left(\xi\right)\right|^{q}d\xi\right)^{\frac{1}{q}}=\left\|\mathcal{E}\right\|_{L^{\infty}\left(U\right)\rightarrow L^{q}\left(B\left(0,R\right)\right)}.$
Theorem 6.
Let $S$ be as above. Suppose that $q>3$ and $0<\nu\leq\frac{1}{2}2^{10}2^{-\frac{3q}{q-3}}$. If $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ holds for all $\varepsilon>0$, then
$Q_{R}^{\left(q\right)}\leq C_{\varepsilon,\nu,q}R^{\varepsilon},\ \ \ \ \ \text{for all }\varepsilon>0\text{ and }R\geq 1.$
Using Theorem 6 together with $\varepsilon$-removal techniques and factorization theory, we can now prove that the Fourier extension conjecture holds in three dimensions if for every $\varepsilon>0$ there is $\nu>0$ such that $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ holds for all $\varepsilon>0$.
Proof of the implication $\left(3\right)\Longrightarrow\left(1\right)$ in Theorem 3.
Statement $\left(3\right)$ of Theorem 3 implies the conclusion of Theorem 6 for all $q>3$, which says that the extension operator $\mathcal{E}$ maps $L^{\infty}\left(\sigma\right)$ to $L^{q}\left(B\left(0,R\right)\right)$ with bound $CR^{\varepsilon}$ for all $q>3$ and $\varepsilon>0$, i.e.
$\left\|\mathcal{E}f\right\|_{L^{q}\left(B\left(0,R\right)\right)}\lesssim R^{\varepsilon}\left\|f\right\|_{L^{\infty}\left(\sigma\right)},\ \ \ \ \ \text{for all }q>3\text{ and }\varepsilon>0.$
By duality, this is equivalent to the restriction inequality,
$\left\|\widehat{f}\mid_{\mathbb{P}^{2}}\right\|_{L^{1}\left(\sigma\right)}\lesssim R^{\varepsilon}\left\|f\right\|_{L^{q^{\prime}}\left(B\left(0,R\right)\right)},\ \ \ \ \ \text{for all }q^{\prime}<\frac{3}{2}\text{ and }\varepsilon>0.$
An immediate consequence of the variant [BoGu, Lemma A1] of Tao’s $\varepsilon$-removal theorem [Tao, Theorem 1.2], is that this inequality implies the global restriction inequality,
$\left\|\widehat{f}\mid_{\mathbb{P}^{2}}\right\|_{L^{1}\left(\sigma\right)}\lesssim\left\|f\right\|_{L^{q^{\prime}}\left(\mathbb{R}^{3}\right)},\ \ \ \ \ \text{for all }q^{\prime}<\frac{3}{2},$
which by duality is the global extension inequality,
$\left\|\mathcal{E}f\right\|_{L^{q}\left(\mathbb{R}^{3}\right)}\lesssim\left\|f\right\|_{L^{\infty}\left(\sigma\right)},\ \ \ \ \ \text{for all }q>3.$
In order to extend the domain $L^{\infty}\left(\sigma\right)$ of $\mathcal{E}$ to the larger space $L^{q}\left(\sigma\right)$, we appeal to Nikishin-Maurey-Pisier factorization theory and interpolation. For example, from [Bus, Corollary 1.4 and Remark 1.5] and Theorem 6, we conclude that,
$\left\|\mathcal{E}g\right\|_{L^{q}\left(\mathbb{R}^{3}\right)}=\left\|\widehat{gd\sigma}\right\|_{L^{q}\left(\mathbb{R}^{3}\right)}\leq C_{q,\gamma}\left\|g\right\|_{L^{q}\left(\sigma\right)},\ \ \ \ \ \text{for all }q>3.$
This completes the proof of Theorem 3 assuming that Theorem 6 holds. ∎
Now we turn to proving the local linear Fourier inequality in Theorem 6 via part of the argument of Bourgain and Guth [BoGu, Section 2], but using the $\nu$-disjoint assumption in (1.2) to simplify the problematic Case 3 of their argument.
2.1. The pigeonholing argument of Bourgain and Guth
Proof of Theorem 6.
We begin the argument exactly as in [BoGu], but with some changes in notation. We let the surface $S$ be a compact smooth piece of the paraboloid $\mathbb{P}^{2}$ given by $z_{3}=\left|z^{\prime}\right|^{2}=z_{1}^{2}+z_{2}^{2}$ in $\mathbb{R}^{3}$, and for $f\in L^{\infty}\left(S\right)$ with $\left\|f\right\|_{L^{\infty}\left(S\right)}=1$, we consider the oscillatory integral $\mathcal{E}f\left(\xi\right)$, which we write as
$Tf\left(\xi\right)$ $\equiv\int_{U}e^{i\phi\left(\xi,y\right)}f\left(y\right)dy=\int_{U}e^{i\left\{\xi_{1}\cdot y_{1}+\xi_{2}\cdot y_{2}+\xi_{3}\left(y_{1}^{2}+y_{2}^{2}\right)\right\}}f\left(y\right)dy$
$=\int_{U}e^{i\xi\cdot\left(y,\left|y\right|^{2}\right)}f\left(y\right)dy=\widehat{f^{\Phi}}\left(\xi\right)=\widehat{\Phi_{\ast}\left[f\left(y\right)dy\right]}\left(\xi\right),\ \ \ \ \ \text{for }\xi\in\mathbb{R}^{3},$
where
$\phi\left(\xi,y\right)=\xi\cdot\Phi\left(y\right)\text{ and }\Phi\left(y\right)\equiv\left(y_{1},y_{2},y_{1}^{2}+y_{2}^{2}\right).$
For $\lambda\geq 1$, let $f=\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}\boldsymbol{1}_{I}f=\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}f_{I}$ and write
$Tf\left(\xi\right)=\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}\int_{S}e^{i\xi\cdot\left(y,\left|y\right|^{2}\right)}f_{I}\left(y\right)dy=\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}e^{i\phi\left(\xi,c_{I}\right)}\int e^{i\left\{\phi\left(\xi,y\right)-\phi\left(\xi,c_{I}\right)\right\}}f_{I}\left(y\right)dy=\sum_{I\in\mathcal{G}_{\lambda}\left[S\right]}e^{i\phi\left(\xi,c_{I}\right)}T_{I}f\left(\xi\right),$
where
(2.2) $T_{I}f\left(\xi\right)$ $\equiv\int e^{i\left\{\phi\left(\xi,y\right)-\phi\left(\xi,c_{I}\right)\right\}}f_{I}\left(y\right)dy=e^{-i\left\{\xi\cdot\Phi\left(c_{I}\right)\right\}}\int e^{i\xi\cdot\Phi\left(y\right)}f_{I}\left(y\right)dy$
$=e^{-i\left\{\xi\cdot\Phi\left(c_{I}\right)\right\}}\widehat{f_{I}^{\Phi}}\left(\xi\right)=\widehat{\tau_{-\Phi\left(c_{I}\right)}f_{I}^{\Phi}}\left(\xi\right),$
where $\tau_{\Phi\left(c_{I}\right)}g\left(z\right)\equiv g\left(z-\Phi\left(c_{I}\right)\right)$ is translation of a function $g$ by the vector $\Phi\left(c_{I}\right)$.
Note that
$\left|\nabla_{\xi}\left\{\xi\cdot\left(y-c_{I},\left|y\right|^{2}-\left|c_{I}\right|^{2}\right)\right\}\right|=\left|\left(y-c_{I},\left|y\right|^{2}-\left|c_{I}\right|^{2}\right)\right|\lesssim\frac{1}{2^{\lambda}},\ \ \ \ \ \text{for }y\in I,$
implies
(2.3) $\nabla_{\xi}T_{I}f\left(\xi\right)$ $=\nabla_{\xi}\int e^{i\xi\cdot\left(y-c_{I},\left|y-c_{I}\right|^{2}\right)}f_{I}\left(y\right)dy=\int\nabla_{\xi}e^{i\xi\cdot\left(y-c_{I},\left|y-c_{I}\right|^{2}\right)}f_{I}\left(y\right)dy$
$=\int ie^{i\xi\cdot\left(y-c_{I},\left|y-c_{I}\right|^{2}\right)}\nabla_{\xi}\left\{\xi\cdot\left(y-c_{I},\left|y\right|^{2}-\left|c_{I}\right|^{2}\right)\right\}f_{I}\left(y\right)dy,$
which implies
(2.4) $\left|\nabla_{\xi}T_{I}f\left(\xi\right)\right|\leq\int\left|\nabla_{\xi}\left\{\xi\cdot\left(y-c_{I},\left|y\right|^{2}-\left|c_{I}\right|^{2}\right)\right\}\right|\left|f_{I}\left(y\right)\right|dy\lesssim\frac{1}{2^{\lambda}}\left\|f_{I}\right\|_{L^{1}\left(U\right)}\lesssim\frac{1}{2^{3\lambda}}\left\|f_{I}\right\|_{L^{\infty}\left(U\right)},$
since $\ell\left(I\right)=\frac{1}{2^{\lambda}}$. We will use the estimates (2.3) and (2.4) in (2.7) below.
Now let $\rho$ be a smooth rapidly decreasing bump function such that $\widehat{\rho}\left(\xi\right)=1$ for $\left|\xi\right|\leq 1$, and set
$\rho_{\lambda}\left(z\right)\equiv\frac{1}{2^{3\lambda}}\rho\left(\frac{z}{2^{\lambda}}\right),\ \ \ \ \ \widehat{\rho_{\lambda}}\left(\xi\right)=\widehat{\rho}\left(2^{\lambda}\xi\right)=1\text{ on }B\left(0,2^{-\lambda}\right)\text{ and }\rho_{\lambda}\left(z\right)\approx\frac{1}{2^{3\lambda}}\text{ on }B\left(0,2^{\lambda}\right).$
Then from (2.2) we obtain
$T_{I}f\left(\xi\right)=T_{I}f\ast\rho_{\lambda}\left(\xi\right)\ ,\ \ \ \ \ \text{for }I\in\mathcal{G}_{\lambda}\left[S\right]\text{ and }\xi\in\mathbb{R}^{3},$
since $\tau_{-\Phi\left(c_{I}\right)}f_{I}^{\Phi}\subset B\left(0,2^{-\lambda}\right)$ and $\widehat{T_{I}f}\left(z\right)=\tau_{-\Phi\left(c_{I}\right)}f_{I}^{\Phi}\left(z\right)$ imply
$\widehat{T_{I}f\ast\rho_{\lambda}}\left(z\right)=\widehat{T_{I}f}\left(z\right)\widehat{\rho_{\lambda}}\left(z\right)=\tau_{-\Phi\left(c_{I}\right)}f_{I}^{\Phi}\left(z\right)\widehat{\rho_{\lambda}}\left(z\right)=\tau_{-\Phi\left(c_{I}\right)}f_{I}^{\Phi}\left(z\right)=\widehat{T_{I}f}\left(z\right).$
Now fix a point $a\in\Gamma_{\lambda}\left(R\right)$, where
$\Gamma_{\lambda}\left(R\right)\equiv 2^{\lambda}\mathbb{Z}^{3}\cap B_{R},\ \ \ \ \ \text{and }B_{R}\equiv B\left(0,R\right),$
and restrict $\xi$ to the ball $B\left(a,2^{\lambda}\right)$. We will write $B_{R}$ as apposed to $B\left(0,R\right)$ above in order to emphasize the nature of the different roles played by $\lambda$ and $R$, namely $R\nearrow\infty$ while $\lambda$ remains a fixed sufficiently large integer to be chosen.
Then for $\xi\in B\left(a,2^{\lambda}\right)$ and $I\in\mathcal{G}_{\lambda}\left[S\right]$ we have
(2.5) $\left|T_{I}f\left(\xi\right)\right|=\left|T_{I}f\ast\rho_{\lambda}\left(\xi\right)\right|=\left|\int_{\mathbb{R}^{3}}T_{I}f\left(z\right)\rho_{\lambda}\left(\xi-z\right)dz\right|$
$\leq\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\left|\rho_{\lambda}\left(\xi-z\right)\right|dz\leq\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\sup_{\omega\in B\left(a,2^{\lambda}\right)}\left|\rho_{\lambda}\left(z-\omega\right)\right|dz=\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz,$
where $\zeta_{\lambda}\left(w\right)\equiv\sup_{\omega-a\in B\left(0,2^{\lambda}\right)}\left|\rho_{\lambda}\left(w-a-\omega\right)\right|$, since $\rho$ can be chosen radial and,
$\sup_{\omega\in B\left(a,2^{\lambda}\right)}\left|\rho_{\lambda}\left(\omega-z\right)\right|$ $=\frac{1}{2^{3\lambda}}\sup_{\omega\in B\left(a,2^{\lambda}\right)}\left|\rho\left(\frac{\left(\omega-a\right)-\left(z-a\right)}{2^{\lambda}}\right)\right|=\frac{1}{2^{3\lambda}}\sup_{\gamma\in B\left(0,1\right)}\left|\rho\left(\frac{z-a}{2^{\lambda}}-\gamma\right)\right|=\zeta_{\lambda}\left(z-a\right),$
$\text{where }\zeta\left(w\right)$ $\equiv\sup_{\left|w-w^{\prime}\right|\leq 1}\left|\rho\left(w^{\prime}\right)\right|.$
Now for $I\in\mathcal{G}_{\lambda}\left[S\right]$ define the right hand side of (2.5) to be
$w_{I}^{a}\left(f\right)\equiv\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz=\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\zeta\left(\frac{z-a}{2^{\lambda}}\right)\frac{dz}{2^{3\lambda}}$
$=\int_{\mathbb{R}^{3}}\left|\widehat{f_{I}^{\Phi}}\left(z\right)\right|\zeta\left(\frac{z-a}{2^{\lambda}}\right)\frac{dz}{2^{3\lambda}}\approx\frac{1}{\left|B\left(a,2^{\lambda}\right)\right|}\int_{B\left(a,2^{\lambda}\right)}\left|\widehat{f_{I}^{\Phi}}\left(z\right)\right|\ ,$
and refer to $w_{I}^{a}\left(f\right)$ as the ‘weight’ of $\widehat{f_{I}^{\Phi}}$ relative to the ball $B\left(a,2^{\lambda}\right)$, which represents that portion of the integral of $\left|\widehat{f_{I}^{\Phi}}\left(z\right)\right|$ that is taken over the ball $B\left(a,2^{\lambda}\right)$. Note that $w_{I}^{a}\left(f\right)\lesssim\left\|\widehat{f_{I}^{\Phi}}\right\|_{L^{\infty}}\lesssim\left\|\mathbf{1}_{I}f\right\|_{L^{1}}\leq\left|I\right|=2^{-2\lambda}$.
Summarizing, we have
(2.6) $\left|T_{I}f\left(\xi\right)\right|\leq\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz=w_{I}^{a}\left(f\right)\ ,\ \ \ \ \ \text{for }\xi\in B\left(a,2^{\lambda}\right).$
and
(2.7) $\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\zeta_{\lambda}\left(z-\xi\right)dz\approx w_{I}^{a}\left(f\right)\text{ },\ \ \ \ \ \text{for }\xi\in B\left(a,2^{\lambda}\right),$
since for $\xi\in B\left(a,2^{\lambda}\right)$ we have $\zeta_{\lambda}\left(\xi-z\right)\approx\zeta_{\lambda}\left(a-z\right)=\zeta_{\lambda}\left(z-a\right)$ by (2.3) and (2.4).
Now set
$w_{\ast}^{a}\left(f\right)\equiv\max_{I\in\mathcal{G}_{\lambda}\left[S\right]}w_{I}^{a}\left(f\right)=\max_{I\in\mathcal{G}_{\lambda}\left[S\right]}\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz,$
and fix $I_{\ast}^{a}$ such that
$w_{I_{\ast}^{a}}^{a}=w_{\ast}^{a}.$
For $1\ll\lambda^{\prime}\ll\lambda$, and $\alpha,\beta,\gamma,\delta\in\mathbb{N}$ chosen appropriately, we will estimate the contributions to the norm $\left\|Tf\right\|_{L^{q}\left(B_{R}\right)}$ in three exhaustive cases in turn. The first case will yield the growth factor $R^{\varepsilon}$, while the next two cases will be absorbed. In fact, we show at the end of the proof that we may take
$\beta=1\text{, \ \ \ \ }\alpha=\delta=\gamma=2\text{, \ \ \ \ }\lambda=2\lambda^{\prime}\text{, \ \ \ \ and }\lambda^{\prime}>\frac{3}{2}\frac{q}{q-3}.$
2.2. Case 1: Separated interaction
Proof continued.
In Case 1 we assume the following property. There exists a triple of squares $I_{0}^{a},J_{0}^{a},K_{0}^{a}\in\mathcal{G}_{\lambda}\left[U\right]$ such that
$w_{I_{0}^{a}},w_{J_{0}^{a}},w_{K_{0}^{a}}$ $>2^{-\alpha\lambda}w_{\ast}^{a}\ ,\ \ \ \ \ \text{and }\left|\mathbf{c}_{I_{0}^{a}}-\mathbf{c}_{J_{0}^{a}}\right|,\left|\mathbf{c}_{J_{0}^{a}}-\mathbf{c}_{K_{0}^{a}}\right|,\left|\mathbf{c}_{K_{0}^{a}}-\mathbf{c}_{I_{0}^{a}}\right|>2^{10}2^{-\beta\lambda}\ ,$
$\framebox{$\begin{array}[c]{ccccc}\mathbf{c}_{I_{0}^{a}}&&&&\\ &\cdot&&&\\ &&\mathbf{c}_{K_{0}^{a}}&\leftrightarrows&\mathbf{c}_{K_{0}^{a}}\\ &&&\cdot&\\ &&&&\mathbf{c}_{J_{0}^{a}}\end{array}$},$
i.e. there exists a $2^{10}2^{-\beta\lambda}$-separated’ triple $I_{0}^{a},J_{0}^{a},K_{0}^{a}$ of squares of side length $2^{-\lambda}$, such that each of $I_{0}^{a}$, $J_{0}^{a}\$and $K_{0}^{a}$ have near maximal weight. In Case 1 we will use the $\nu$-disjoint trilinear estimate in Theorem 3 with $\nu=2^{10}2^{-\beta\lambda}$, and $U_{I_{0}^{a}}$ equal to a cube of side length $\mathop{\rm dist}\left(I_{0}^{a},J_{0}^{a}\cup K_{0}^{a}\right)$, and similarly for $U_{J_{0}^{a}}$ and $U_{K_{0}^{a}}$. For $\xi\in B\left(a,2^{\lambda}\right)$ we throw away the unimodular function $e^{-i\Phi\left(c_{I}\right)\cdot\xi}$, and using (2.6), we estimate that for $\xi\in B\left(a,2^{\lambda}\right)$,
(2.8) $\left|Tf\left(\xi\right)\right|=\left|\sum_{L\in\mathcal{G}_{\lambda}\left[S\right]}e^{i\phi\left(\xi,c_{I}\right)}T_{L}f\left(\xi\right)\right|\leq\sum_{L\in\mathcal{G}_{\lambda}\left[S\right]}\left|T_{L}f\left(\xi\right)\right|\lesssim\sum_{L\in\mathcal{G}_{\lambda}\left[S\right]}w_{L}^{a}<2^{2\lambda}w_{\ast}^{a}<2^{\left(2+\alpha\right)\lambda}\left(w_{I_{0}^{a}}w_{J_{0}^{a}}w_{K_{0}^{a}}\right)^{\frac{1}{3}},$
since the fixed triple $\left(I_{0}^{a},J_{0}^{a},K_{0}^{a}\right)$ satisfies the near maximal weight condition in Case 1:
$w_{\ast}^{a}<\min\left\{2^{\alpha\lambda}w_{I_{0}^{a}}^{a},2^{\alpha\lambda}w_{J_{0}^{a}}^{a},2^{\alpha\lambda}w_{K_{0}^{a}}^{a}\right\}\leq 2^{\alpha\lambda}\left(w_{I_{0}^{a}}^{a}\right)^{\frac{1}{3}}\left(w_{J_{0}^{a}}^{a}\right)^{\frac{1}{3}}\left(w_{K_{0}^{a}}^{a}\right)^{\frac{1}{3}}.$
Let $\nu=2^{10}2^{-\beta\lambda}$. Then for $q>3$ and $\xi\in B\left(a,2^{\lambda}\right)$, we have from boundedness of $Tf$, and (2.8) and (2.7), followed by Hölder’s inequality, that
$\left|Tf\left(\xi\right)\right|^{q}\lesssim 2^{q\left(2+\alpha\right)\lambda}\left(w_{I_{0}^{a}}^{a}w_{J_{0}^{a}}^{a}w_{K_{0}^{a}}^{a}\right)^{\frac{q}{3}}$
$\approx 2^{q\left(2+\alpha\right)\lambda}\left(\int_{\mathbb{R}^{3}}\left|T_{I_{0}^{a}}f\left(z_{1}\right)\right|\zeta_{\lambda}\left(z_{1}-a\right)dz_{1}\right)^{\frac{q}{3}}\left(\int_{\mathbb{R}^{3}}\left|T_{J_{0}^{a}}f\left(z_{2}\right)\right|\zeta_{\lambda}\left(z_{2}-a\right)dz_{2}\right)^{\frac{q}{3}}\left(\int_{\mathbb{R}^{3}}\left|T_{K_{0}^{a}}f\left(z_{3}\right)\right|\zeta_{\lambda}\left(z_{3}-a\right)dz_{3}\right)^{\frac{q}{3}}$
$\lesssim 2^{q\left(2+\alpha\right)\lambda}\int_{\mathbb{R}^{3}}\int_{\mathbb{R}^{3}}\int_{\mathbb{R}^{3}}\left|T_{I_{0}^{a}}f\left(\xi-z_{1}\right)T_{J_{0}^{a}}f\left(\xi-z_{2}\right)T_{K_{0}^{a}}f\left(\xi-z_{3}\right)\right|^{\frac{q}{3}}\zeta_{\lambda}\left(z_{1}\right)\zeta_{\lambda}\left(z_{2}\right)\zeta_{\lambda}\left(z_{3}\right)dz_{1}dz_{2}dz_{3}$
$\lesssim 2^{q\left(2+\alpha\right)\lambda}\sum_{\left(I,J,K\right)\in\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]}\int_{\mathbb{R}^{9}}\left|T_{I}f\left(\xi-z_{1}\right)T_{J}f\left(\xi-z_{2}\right)T_{K}f\left(\xi-z_{3}\right)\right|^{\frac{q}{3}}d\mu_{\lambda}\left(z_{1},z_{2},z_{3}\right)\ ,$
where
[…]
and
$d\mu_{\lambda}\left(z_{1},z_{2},z_{3}\right)\equiv\zeta_{\lambda}\left(z_{1}\right)\zeta_{\lambda}\left(z_{2}\right)\zeta_{\lambda}\left(z_{3}\right)dz_{1}dz_{2}dz_{3}$
is a bounded multiple of a probability measure.
Then we have
$\int_{B\left(a,2^{\lambda}\right)}\left|Tf\left(\xi\right)\right|^{q}d\xi$
$\lesssim 2^{q\left(2+\alpha\right)\lambda}\int_{B\left(a,2^{\lambda}\right)}\sum_{\left(I,J,K\right)\in\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]}\int_{\mathbb{R}^{9}}\left|T_{I}f\left(\xi-z_{1}\right)T_{J}f\left(\xi-z_{2}\right)T_{K}f\left(\xi-z_{3}\right)\right|^{\frac{q}{3}}d\mu_{\lambda}\left(z_{1},z_{2},z_{3}\right)d\xi$
$=2^{q\left(2+\alpha\right)\lambda}\sum_{\left(I,J,K\right)\in\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]}\int_{\mathbb{R}^{9}}\left\{\int_{B\left(a,2^{\lambda}\right)}\left|T_{I}f\left(\xi-z_{1}\right)T_{J}f\left(\xi-z_{2}\right)T_{K}f\left(\xi-z_{3}\right)\right|^{\frac{q}{3}}d\xi\right\}d\mu_{\lambda}\left(z_{1},z_{2},z_{3}\right).$
Now consider those $a\in\Gamma_{\lambda}\left(R\right)$ for which Case 1 is in effect for the ball $B\left(a,2^{\lambda}\right)$ and denote by $\Gamma_{\lambda}\left(\text{{Case 1}}\right)$ the union of all the balls $B\left(a,2^{\lambda}\right)$ for which $a$ is in Case 1. Summing over points $a\in\Gamma_{\lambda}\left(R\right)$ such that Case 1 is in effect for the ball $B\left(a,2^{\lambda}\right)$, we obtain
$\sum_{a\in\Gamma_{\lambda}\left(\text{{Case 1}}\right)}\int_{B\left(a,2^{\lambda}\right)}\left|Tf\left(\xi\right)\right|^{q}d\xi$
$\lesssim 2^{q\left(2+\alpha\right)\lambda}\sum_{\left(I,J,K\right)\in\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]}\int_{\mathbb{R}^{9}}\left\{\sum_{a\in\Gamma_{\lambda}\left(\text{{Case 1}}\right)}\int_{B\left(a,2^{\lambda}\right)}\left|T_{I}f\left(\xi-z_{1}\right)T_{J}f\left(\xi-z_{2}\right)T_{K}f\left(\xi-z_{3}\right)\right|^{\frac{q}{3}}d\xi\right\}d\mu_{\lambda}\left(z_{1},z_{2},z_{3}\right)$
$\lesssim 2^{q\left(2+\alpha\right)\lambda}\sum_{\left(I,J,K\right)\in\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]}\int_{\mathbb{R}^{9}}\left\{\int_{B_{R}}\left|T_{I}f\left(\xi-z_{1}\right)T_{J}f\left(\xi-z_{2}\right)T_{K}f\left(\xi-z_{3}\right)\right|^{\frac{q}{3}}d\xi\right\}d\mu_{\lambda}\left(z_{1},z_{2},z_{3}\right)$
$\leq 2^{q\left(2+\alpha\right)\lambda}\left(\#\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]\right)\int_{\mathbb{R}^{9}}\left\{C_{\varepsilon}^{q}R^{q\varepsilon}\right\}d\mu_{\lambda}\left(z_{1},z_{2}...,z_{N}\right)\lesssim C_{\varepsilon}^{q}2^{q\left(2+\alpha\right)\lambda}2^{6\lambda}R^{q\varepsilon},$
upon appealing to the $\nu$-disjoint trilinear assumption $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{p}\rightarrow L^{\frac{p}{3}}\right)$ in part (3) of Theorem 3 with $\nu=2^{10}2^{-\lambda}$, and
$f_{1}=\widetilde{\mathsf{M}_{z_{1}}}f_{I},\ \ \ f_{2}=\widetilde{\mathsf{M}_{z_{1}}}f_{J},\ \ \ \ f_{3}=\widetilde{\mathsf{M}_{z_{1}}}f_{K},$
where $\widetilde{\mathsf{M}_{z}}\left(x\right)=e^{i\left\langle z,\Phi\left(x\right)\right\rangle}$. Indeed, if $\mathsf{M}_{z}\left(w\right)=e^{i\left\langle z,w\right\rangle}$ then $\mathsf{M}_{z}\Phi_{\ast}=\Phi_{\ast}\widetilde{\mathsf{M}_{z}}$ since for $\varphi$ continuous, and with the pushforward and pullback operators $\Phi_{\ast},\Phi^{\ast}$, we have
$\left\langle\mathsf{M}_{z}\Phi_{\ast}g,\varphi\right\rangle$ $=\int\left\{e^{i\left\langle z,w\right\rangle}\Phi_{\ast}g\left(w\right)\right\}\varphi\left(w\right)dw=\int\left\{e^{i\left\langle z,w\right\rangle}\varphi\left(w\right)\right\}\Phi_{\ast}g\left(w\right)dw=\int\left\{e^{i\left\langle z,\Phi\left(x\right)\right\rangle}\varphi\left(\Phi\left(x\right)\right)\right\}g\left(x\right)dx$
$=\int\left\{e^{i\left\langle z,\Phi\left(x\right)\right\rangle}g\left(x\right)\right\}\Phi^{\ast}\varphi\left(x\right)dx=\left\langle\widetilde{\mathsf{M}_{z}}g,\Phi^{\ast}\varphi\right\rangle=\left\langle\Phi_{\ast}\widetilde{\mathsf{M}_{z}}g,\varphi\right\rangle.$
Thus,
$\left|T_{I}f\left(\xi-z_{1}\right)\right|$ $=\widehat{\Phi_{\ast}f_{I}}\left(\xi-z_{1}\right)=\widehat{\mathsf{M}_{z_{1}}\Phi_{\ast}f_{I}}\left(\xi\right)=\widehat{\Phi_{\ast}\widetilde{\mathsf{M}_{z_{1}}}f_{I}}\left(\xi\right)=\mathcal{E}f_{1}\left(\xi\right),$
$\text{and }\left|T_{J}f\left(\xi-z_{2}\right)\right|$ $=\mathcal{E}f_{1}\left(\xi\right)\text{ and }\left|T_{K}f\left(\xi-z_{1}\right)\right|=\mathcal{E}f_{3}\left(\xi\right)\text{,}$
and so
$\int_{B_{R}}\left|T_{I}f\left(\xi-z_{1}\right)T_{J}f\left(\xi-z_{2}\right)T_{K}f\left(\xi-z_{3}\right)\right|^{\frac{q}{3}}d\xi=\int_{B_{R}}\left|\mathcal{E}f_{1}\left(\xi\right)\mathcal{E}f_{2}\left(\xi\right)\mathcal{E}f_{3}\left(\xi\right)\right|^{\frac{q}{3}}d\xi$
$=\left\|\prod_{j=1}^{3}\mathcal{E}_{j}f_{j}\right\|_{L^{{}^{\frac{q}{3}}}\left(B_{R}\right)}^{\frac{q}{3}}\leq\left(C_{\varepsilon,\nu,q}R^{\varepsilon}\right)^{q}\prod_{j=1}^{3}\left\|f_{j}\right\|_{L^{\infty}}^{\frac{q}{3}}=\left(C_{\varepsilon,\nu,q}R^{\varepsilon}\right)^{q},$
since the triple $\left(I,J,K\right)$ is $\nu$-separated as in (1.3), and since $\left|f_{j}\right|\leq 1$. As a consequence we have
$\sum_{\left(I,J,K\right)\in\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]}\int_{B_{R}}\left|T_{I}f\left(\xi\right)T_{J}f\left(\xi\right)T_{K}f\left(\xi\right)\right|^{\frac{q}{3}}d\xi$
$\leq\sum_{\left(I,J,K\right)\in\mathcal{G}_{\lambda}^{\nu-\mathop{\rm separated}}\left[U\right]}\left(C_{\varepsilon,\nu,q}R^{\varepsilon}\right)^{q}\lesssim 2^{6\lambda}\left(C_{\varepsilon,\nu,q}R^{\varepsilon}\right)^{q}=\left(C_{\varepsilon,\nu,q}\right)^{q}2^{6\lambda}R^{q\varepsilon}.$
Altogether then, we have proved that
$\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 1}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}\lesssim C_{\varepsilon,\nu,q}2^{\left(\frac{6}{q}+2+\alpha\right)\lambda}R^{\varepsilon},$
where $\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 1}}\right)}$ indicates the union of those balls $B\left(a,2^{\lambda}\right)$ for which Case 1 holds. ∎
2.3. Case 2: Clustered interaction
Proof continued.
In Case 2 we assume the following property. If $\left|c_{I}-c_{I_{\ast}^{a}}\right|>2^{-\gamma\lambda^{\prime}}$, then $w_{I}^{a}\leq 2^{-\delta\lambda}w_{\ast}^{a}$. In other words, if $I$ is sufficiently far from $I_{\ast}^{a}$, then $w_{I}^{a}$ is much smaller than $w_{\ast}^{a}$, i.e.
$\mathop{\rm dist}\left(I,I_{\ast}^{a}\right)$ $>2^{-\gamma\lambda^{\prime}}\Longrightarrow\int_{\mathbb{R}^{3}}\left|T_{I}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz\leq 2^{-\delta\lambda}\int_{\mathbb{R}^{3}}\left|T_{I_{\ast}^{a}}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz,$
$\framebox{$\begin{array}[c]{ccccc}\mathbf{c}_{I}&&&&\\ &\ast&&&\\ &&\ast&&\\ &&&\ast&\\ &&&&\mathbf{c}_{I_{\ast}^{a}}\end{array}$}.$
In this case we will use rescaling and recursion as in [TaVaVe].
Let $\xi\in B\left(a,2^{\lambda}\right)$ for some $a\in\Gamma_{\lambda}\left(R\right)$. Using that $\left|c_{I}-c_{I_{\ast}^{a}}\right|>2^{-\gamma\lambda^{\prime}}$ implies $w_{I}^{a}\leq 2^{-\delta\lambda}w_{\ast}^{a}$ in this case, we have with $\left|\cdot\right|_{\mathop{\rm square}}$ denoting the ‘square’ norm in $\mathbb{R}^{3}$,
$\left|Tf\left(\xi\right)\right|$ $=\left|\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}\int_{I}e^{i\phi\left(\xi,y\right)}f\left(y\right)dy\right|$
$\leq\left|\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]:\ \left|c_{I}-c_{I_{\ast}^{a}}\right|_{\mathop{\rm square}}\leq 2^{-\lambda^{\prime}}}\int_{I}e^{i\phi\left(\xi,y\right)}f\left(y\right)dy\right|+\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]:\ \left|c_{I}-c_{I_{\ast}^{a}}\right|>2^{-\lambda^{\prime}}}\left|T_{I}f\left(\xi\right)\right|$
$\leq 10\max_{K\in\mathcal{G}_{\lambda^{\prime}}\left[U\right]}\left|\int_{K}e^{i\phi\left(\xi,y\right)}f\left(y\right)dy\right|+\sum_{\left|c_{I}-c_{I_{\ast}^{a}}\right|>2^{-\lambda^{\prime}}}w_{I}^{a}$
$\leq 10\max_{K\in\mathcal{G}_{\lambda^{\prime}}\left[U\right]}\left|T_{K}f\left(\xi\right)\right|+\left(\#\mathcal{G}_{\lambda}\left[U\right]\right)2^{-\delta\lambda}w_{\ast}^{a}\leq 10\max_{K\in\mathcal{G}_{\lambda^{\prime}}\left[U\right]}\left|T_{K}f\left(\xi\right)\right|+2^{\left(2-\delta\right)\lambda}w_{\ast}^{a}\ ,$
since if $K_{\ast}^{a}\in\mathcal{G}_{\lambda^{\prime}}\left[S\right]$ contains $I_{\ast}^{a}$, then we decompose $f=f_{K_{\ast}^{a}}+\sum_{I\in\mathcal{G}_{\lambda}\left[S\right]:\ I\cap K_{\ast}=\emptyset}f_{I}$, and without loss of generality we may also assume $\left|c_{I}-c_{I_{\ast}^{a}}\right|\gtrsim 2^{-\gamma\lambda^{\prime}}$. Now
$\int\left|T_{I}f\left(\xi-z\right)\right|\zeta_{\lambda}^{a}\left(z\right)dz$ $\leq\left(\int\left|T_{I}f\left(\xi-z\right)\right|^{q}\zeta_{\lambda}\left(z-a\right)dz\right)^{\frac{1}{q}}\left(\int\zeta_{\lambda}\left(z-a\right)dz\right)^{\frac{1}{q^{\prime}}}$
$\lesssim\left(\int\left|T_{I}f\left(z-a\right)\right|^{q}\zeta_{\lambda}\left(z\right)dz\right)^{\frac{1}{q}},$
and so for $\xi\in B\left(a,2^{\lambda}\right)$,
$\left|Tf\left(\xi\right)\right|^{q}$ $\leq C\sum_{K\in\mathcal{G}_{\lambda^{\prime}}\left[U\right]}\left|T_{K}f\left(\xi\right)\right|^{q}+C2^{\left(2-\delta\right)\lambda q}\int\left|T_{I_{\ast}^{a}}f\left(z\right)\right|^{q}\zeta_{\lambda}\left(z-a\right)dz$
$\leq C\sum_{K\in\mathcal{G}_{\lambda^{\prime}}\left[U\right]}\left|T_{K}f\left(\xi\right)\right|^{q}+C2^{\left(2-\delta\right)\lambda q}\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}\int\left|T_{I}f\left(z\right)\right|^{q}\zeta_{\lambda}\left(z-a\right)dz,$
where we have added in all $K\in\mathcal{G}_{\lambda^{\prime}}\left[S\right]$ rather than just $K_{\ast}^{a}$, and all $I\in\mathcal{G}_{\lambda}\left[S\right]$ rather than just $I_{\ast}^{a}$.
Summing over $a\in\Gamma_{\lambda}\left(R\right)$, we see that the corresponding contribution over $B_{R}$ is at most
$\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 2}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}^{q}$ $\equiv\sum_{a\in\Gamma_{\lambda}\left(R\right)}\left\{C\sum_{K\in\mathcal{G}_{\lambda^{\prime}}\left[U\right]}\int_{B\left(a,2^{s}\right)}\left|T_{K}f\left(\xi\right)\right|^{q}d\xi+C2^{\left(2-\delta\right)\lambda q}\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}\int\left|T_{I}f\left(z\right)\right|^{q}\zeta_{s}^{a}\left(z\right)dz\right\}$
$\lesssim C\sum_{K\in\mathcal{G}_{\lambda^{\prime}}\left[U\right]}\int_{B_{R}}\left|T_{K}f\left(\xi\right)\right|^{q}d\xi+C2^{-3\lambda}2^{\left(2-\delta\right)\lambda q}\sum_{I\in\mathcal{G}_{\lambda}\left[U\right]}\int_{B_{R}}\left|T_{I}f\left(\xi\right)\right|^{q}d\xi,$
since $\sum_{a\in\Gamma_{\lambda}\left(R\right)}\zeta_{\lambda}\left(z-a\right)\lesssim 2^{-3\lambda}\mathbf{1}_{B_{R}}\left(z\right)+\mathop{\rm rapid}\mathop{\rm decay}$.
At this point we follow [BoGu] in using parabolic rescaling, as introduced in Tao, Vargas and Vega [TaVaVe], on the integral
$\mathop{\rm Int}_{\rho}\left(\xi\right)\equiv\int_{\left|y-\overline{y}\right|<\rho}e^{i\phi\left(\xi,y\right)}f\left(y\right)dy=\int_{\left|y-\overline{y}\right|<\rho}e^{i\left[\xi_{1}y_{1}+\xi_{2}y_{2}+\xi_{2}\left(y_{1}^{2}+y_{2}^{2}\right)\right]}f\left(y\right)dy,\ \ \ \ \ \text{for }0<\rho<1,$
to obtain
$\left|\mathop{\rm Int}_{\rho}\left(\xi\right)\right|\overset{y=\overline{y}+y^{\prime}}{=}\left|\int_{\left|y^{\prime}\right|<\rho}e^{i\left[\xi_{1}\left(\overline{y}_{1}+y_{1}^{\prime}\right)+\xi_{2}\left(\overline{y}_{2}+y_{2}^{\prime}\right)+\xi_{3}\left(\left(\overline{y}_{1}+y_{1}^{\prime}\right)^{2}+\left(\overline{y}_{2}+y_{2}^{\prime}\right)^{2}\right)\right]}f\left(\overline{y}+y^{\prime}\right)dy^{\prime}\right|$
$=\left|\int_{\left|y^{\prime}\right|<\rho}e^{i\left[\left(\xi_{1}+2\overline{y_{1}}\xi_{3}\right)y_{1}^{\prime}+\left(\xi_{2}+2\overline{y_{2}}\xi_{3}\right)y_{2}^{\prime}+\xi_{3}\left|y^{\prime}\right|^{2}\right]}f\left(\overline{y}+y^{\prime}\right)dy^{\prime}\right|.$
Thus we conclude that
(2.10) $\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left\|\mathop{\rm Int}_{\rho}\right\|_{L^{q}\left(B_{R}\right)}=\left(\int_{B_{R}}\left|\mathop{\rm Int}_{\rho}\left(\xi\right)\right|^{q}d\xi\right)^{\frac{1}{q}}$
$=\left(\int_{B_{R}}\left|\int_{\left|y^{\prime}\right|<\rho}e^{i\left[\left(\xi_{1}+2\overline{y_{1}}\xi_{3}\right)y_{1}^{\prime}+\left(\xi_{2}+2\overline{y_{2}}\xi_{3}\right)y_{2}^{\prime}+\xi_{3}\left|y^{\prime}\right|^{2}\right]}f\left(\overline{y}+y^{\prime}\right)dy^{\prime}\right|^{q}d\xi\right)^{\frac{1}{q}}$
$=\left(\int_{B_{R}}\left|\int_{\left|y^{\prime}\right|<\rho}e^{i\left[\left(\rho\xi_{1}+2\overline{\frac{y_{1}}{\rho}}\rho^{2}\xi_{3}\right)\frac{y_{1}^{\prime}}{\rho}+\left(\rho\xi_{2}+2\overline{\frac{y_{2}}{\rho}}\rho^{2}\xi_{3}\right)y_{2}^{\prime}+\rho^{2}\xi_{3}\left|\frac{y^{\prime}}{\rho}\right|^{2}\right]}f\left(\overline{y}+y^{\prime}\right)\rho^{2}d\left(\frac{y^{\prime}}{\rho}\right)\right|^{q}\frac{d\left(\rho\xi^{\prime}\right)d\left(\rho^{2}\xi_{3}\right)}{\rho^{4}}\right)^{\frac{1}{q}}$
$=\rho^{2}\rho^{-\frac{4}{q}}\left(\int_{B_{\rho R}}\left|\int_{\left|y^{\prime}\right|<1}e^{i\left[\left(\xi_{1}+2\overline{y_{1}}\xi_{3}\right)y_{1}^{\prime}+\left(\xi_{2}+2\overline{y_{2}}\xi_{3}\right)y_{2}^{\prime}+\xi_{3}\left|y^{\prime}\right|^{2}\right]}f\left(\rho\left(\overline{y}+y^{\prime}\right)\right)dy^{\prime}\right|^{q}d\xi^{\prime}d\xi_{3}\right)^{\frac{1}{q}}\leq C\rho^{2}\rho^{-\frac{4}{q}}Q_{\rho R}^{\left(q\right)}\ ,$
where $Q_{\rho R}^{\left(q\right)}\leq Q_{R}^{\left(q\right)}$ is defined in (2.1), since
[…]
the paraboloid is invariant under parabolic rescaling.
Note that the factor $\rho^{2}$ arises from $\left|y^{\prime}\right|<\rho$, and that the factor $\rho^{-\frac{4}{q}}Q_{\rho R}^{\left(q\right)}$ arises from parabolic rescaling. These features remain in play for an arbitrary quadratic surface of positive Gaussian curvature.
Thus using (2.10), first with $\rho=2^{-\lambda^{\prime}}$ and then with $\rho=2^{-\lambda}$, we obtain
$\left\|T_{K}f\right\|_{L^{q}\left(B_{R}\right)}\lesssim 2^{-\left(2-\frac{4}{q}\right)\lambda^{\prime}}Q_{2^{-\lambda^{\prime}}R}^{\left(q\right)}\text{ and }\left\|T_{I}f\right\|_{L^{q}\left(B_{R}\right)}\lesssim 2^{-\left(2-\frac{4}{q}\right)\lambda}Q_{2^{-\lambda}R}^{\left(q\right)}\ ,$
and together with (2.9), we obtain that the contribution $\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 2}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}$ to the norm $\left\|Tf\right\|_{L^{q}\left(B_{R}\right)}$ satisfies:
(2.12) $\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 2}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}\leq C\left(\#\mathcal{G}_{\lambda^{\prime}}\left[S\right]\right)^{\frac{1}{q}}\left(2^{-\lambda^{\prime}}\right)^{2-\frac{4}{q}}Q_{2^{-\lambda^{\prime}}R}^{\left(q\right)}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +C2^{-\left(\delta-2+\frac{3}{q}\right)\lambda}\left(\#\mathcal{G}_{\lambda}\left[S\right]\right)^{\frac{1}{q}}\left(2^{-\lambda}\right)^{2-\frac{4}{q}}Q_{2^{-\lambda}R}^{\left(q\right)}$
$=C2^{\left(\frac{6}{q}-2\right)\lambda^{\prime}}Q_{2^{-\lambda^{\prime}}R}^{\left(q\right)}+C2^{\left(\frac{3}{q}-\delta\right)\lambda}Q_{2^{-\lambda}R}^{\left(q\right)}\ ,$
since
$\left(\#\mathcal{G}_{\lambda^{\prime}}\left[S\right]\right)^{\frac{1}{q}}\left(2^{-\lambda^{\prime}}\right)^{2}\left(2^{-\lambda^{\prime}}\right)^{-\frac{4}{q}}=2^{\frac{2}{q}\lambda^{\prime}}2^{-2\lambda^{\prime}}2^{\frac{4}{q}\lambda^{\prime}}=2^{\left(\frac{6}{q}-2\right)\lambda^{\prime}},$
and
$2^{-\left(\delta-2+\frac{3}{q}\right)\lambda}\left(\#\mathcal{G}_{\lambda}\left[S\right]\right)^{\frac{1}{q}}\left(2^{-\lambda}\right)^{2-\frac{4}{q}}=2^{-\frac{3}{q}\lambda}2^{\left(2-\delta\right)\lambda}2^{\frac{2}{q}\lambda}2^{-2\lambda}2^{\frac{4}{q}\lambda}=2^{\left(\frac{3}{q}-\delta\right)\lambda}\ .$
2.4. Case 3: Dipole interaction
Proof continued.
In Case 3 we assume the negation of both Case 1 and Case 2. The failure of clustered interaction implies that there also exists $I_{\ast\ast}^{a}$ with $w_{I_{\ast\ast}^{a}}>2^{-\delta\lambda}w_{\ast}^{a}$ and $\left|c_{I_{\ast\ast}^{a}}-c_{I_{\ast}^{a}}\right|>2^{-\gamma\lambda^{\prime}}$, i.e.
$\int_{\mathbb{R}^{3}}\left|T_{I_{\ast\ast}^{a}}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz$ $=w_{I_{\ast\ast}^{a}}>2^{-\delta\lambda}w_{\ast}^{a}=2^{-\delta\lambda}\int_{\mathbb{R}^{3}}\left|T_{I_{\ast}^{a}}f\left(z\right)\right|\zeta_{\lambda}\left(z-a\right)dz\ ,$
$\mathop{\rm dist}\left(I_{\ast\ast},I_{\ast}\right)>2^{-\gamma\lambda^{\prime}}\ ,$
The simultaneous failure of separated interaction further implies that
(2.13) $w_{I}^{a}$ $\leq 2^{-\alpha\lambda}w_{\ast}^{a}\text{ if }\mathop{\rm dist}\left(c_{I},I_{\ast}^{a}\cup I_{\ast\ast}^{a}\right)>2^{10}2^{-\beta\lambda},$
(2.20) $\framebox{$\begin{array}[c]{ccccccc}\mathbf{c}_{I_{\ast\ast}^{a}}&&&&&&\\ &\ast&&&&&\\ &&\ast&&&&\\ &&&\ast&&&\\ &&&&\mathbf{c}_{I_{\ast}^{a}}&&\\ &&&&&&\mathbf{c}_{I}^{a}\end{array}$}$
In this case we will again use parabolic rescaling since the squares $I$ with near maximal weight, i.e. $2^{-\alpha\lambda}w_{\ast}^{a}$, are clustered within distance $2^{10}2^{-\beta\lambda}$ of the squares $I_{\ast}^{a}$ and $I_{\ast\ast}^{a}$. Indeed, arguing as in (2.9) and (2.12) above, we then have
$\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 3}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}$
$\lesssim\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 3}}\right)}\sum_{\begin{subarray}{c}I:\ \mathop{\rm dist}\left(c_{I},I_{\ast}^{a}\right)\leq 2^{10}2^{-\lambda^{\prime}}\\ \text{or }\mathop{\rm dist}\left(c_{I},I_{\ast}^{a}\cup I_{\ast\ast}^{a}\right)>2^{10}2^{-\lambda^{\prime}}\end{subarray}}T_{I}f\right\|_{L^{q}\left(B_{R}\right)}+\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 3}}\right)}\sum_{I:\ \mathop{\rm dist}\left(c_{I},I_{\ast\ast}^{a}\right)\leq 2^{10}2^{-\lambda^{\prime}}}T_{I}f\right\|_{L^{q}\left(B_{R}\right)}$
$\lesssim C2^{\left(\frac{6}{q}-2\right)\lambda^{\prime}}Q_{2^{-\lambda^{\prime}}R}^{\left(q\right)}+C2^{\left(\frac{3}{q}-\delta\right)\lambda}Q_{2^{-\lambda}R}^{\left(q\right)}\ .$
Note that we needed only to consider the squares satisfying $\mathop{\rm dist}\left(c_{I},I_{\ast}^{a}\cup I_{\ast\ast}^{a}\right)>2^{10}2^{-\beta\lambda}$ in (2.9), for just one of the dipoles $I_{\ast}^{a}$ or $I_{\ast\ast}^{a}$. ∎
2.5. Completing the proof
Proof continued.
So far we have shown that if $\lambda^{\prime}=\frac{\beta}{\gamma}\lambda$ and
$0<\beta<\gamma\ \text{and }0<\alpha=\delta\leq 3,$
then
$\left\|Tf\right\|_{L^{q}\left(B_{R}\right)}$ $\leq\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 1}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}+\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 2}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}+\left\|\mathbf{1}_{\Gamma_{\lambda}\left(\text{{Case 3}}\right)}Tf\right\|_{L^{q}\left(B_{R}\right)}$
$\lesssim C_{\varepsilon,\nu,q}2^{\left(\frac{6}{q}+2+\alpha\right)\lambda}R^{\varepsilon}+C2^{\left(\frac{6}{q}-2\right)\lambda^{\prime}}Q_{2^{-\lambda^{\prime}}R}^{\left(q\right)}+C2^{\left(\frac{3}{q}-\delta\right)\lambda}Q_{2^{-\lambda}R}^{\left(q\right)}+\left(2^{-\beta\lambda}\right)^{2-\frac{4}{q}}Q_{2^{-\beta\lambda}R}^{\left(q\right)}$
$\lesssim C_{\varepsilon,\nu,q}2^{\left(\frac{6}{q}+2+\alpha\right)\lambda}R^{\varepsilon}+\left[2^{-\frac{2}{q}\left(q-3\right)\lambda^{\prime}}+2^{-\left(\delta-\frac{3}{q}\right)\lambda}+2^{-\beta\left(2-\frac{4}{q}\right)\lambda}\right]Q_{R}^{\left(q\right)}$
$\lesssim C_{\varepsilon,\nu,q}2^{\left(\frac{6}{q}+2+\alpha\right)\lambda}R^{\varepsilon}+\frac{1}{2}Q_{R}^{\left(q\right)}\ ,$
if $q>3$ and both $\lambda$ and $\lambda^{\prime}$ are sufficiently large, namely
(2.21) $2^{-\left(q-3\right)\frac{2}{q}\frac{\beta}{\gamma}\lambda}+2^{-\left(\delta-\frac{3}{q}\right)\lambda}+2^{-\beta\left(2-\frac{4}{q}\right)\lambda}<\frac{1}{2}.$
If we take $\beta=1$ and $\alpha=\delta=\gamma=2$, then (2.21) becomes,
$2^{-\left(1-\frac{3}{q}\right)\lambda}+2^{-\left(2-\frac{3}{q}\right)\lambda}+2^{-\left(2-\frac{4}{q}\right)\lambda}<\frac{1}{2},$
which is satisfied if
$\left(1-\frac{3}{q}\right)\lambda,\left(2-\frac{4}{q}\right)\lambda\geq 3\text{, in particular if }\lambda=\frac{3q}{q-3}.$
Thus
$Q_{R}^{\left(q\right)}=\sup_{\left\|f\right\|_{L^{\infty}}\leq 1}\left\|Tf\right\|_{L^{q}\left(B_{R}\right)}\lesssim C_{\varepsilon,\nu,q}2^{\left(\frac{6}{q}+4\right)\lambda}R^{\varepsilon}+\frac{1}{2}Q_{R}^{\left(q\right)},$
and absorption now yields the inequality,
$Q_{R}^{\left(q\right)}\leq 2C_{\varepsilon,\nu,q}2^{\left(\frac{6}{q}+4\right)\lambda}R^{\varepsilon}\leq 2C_{\varepsilon,\nu,q}2^{\left(\frac{6}{q}+4\right)\frac{3q}{q-3}}R^{\varepsilon}\leq 2C_{\varepsilon,\nu,q}2^{18\frac{q}{q-3}}R^{\varepsilon},\ \ \ \ \ \text{for all }R\geq 1.$
Thus the disjoint constant $\nu=2^{10}2^{-\beta\lambda}$ is given by
(2.22) $\nu=2^{10}2^{-\frac{3q}{q-3}},$
which depends only on how much larger $q$ is than $3$. This completes the proof of Theorem 6. ∎
3. A smooth Alpert characterization
First we recall the construction from [Saw7] of smooth Alpert projections $\left\{\bigtriangleup_{Q;\kappa}\right\}_{Q\in\mathcal{D}}$ and corresponding wavelets $\left\{h_{Q;\kappa}^{a}\right\}_{Q\in\mathcal{D},\ a\in\Gamma_{n}}$ of order $\kappa$ in $n$-dimensional space $\mathbb{R}^{n}$, giving fairly complete definitions and statements. In fact, $\left\{h_{Q;\kappa}^{a}\right\}_{a\in\Gamma}$ is an orthonormal basis for the finite dimensional vector subspace of $L^{2}$ that consists of linear combinations of the indicators of the children $\mathfrak{C}\left(Q\right)$ of $Q$ multiplied by polynomials of degree at most $\kappa-1$, and such that the linear combinations have vanishing moments on the cube $Q$ up to order $\kappa-1$:
$L_{Q;k}^{2}\left(\mu\right)\equiv\left\{f=\mathop{\displaystyle\sum}\limits_{Q^{\prime}\in\mathfrak{C}\left(Q\right)}\mathbf{1}_{Q^{\prime}}p_{Q^{\prime};k}\left(x\right):\int_{Q}f\left(x\right)x_{i}^{\ell}d\mu\left(x\right)=0,\ \ \ \text{for }0\leq\ell\leq k-1\text{ and }1\leq i\leq n\right\},$
where $p_{Q^{\prime};k}\left(x\right)=\sum_{\alpha\in\mathbb{Z}_{+}^{n}:\left|\alpha\right|\leq k-1\ }a_{Q^{\prime};\alpha}x^{\alpha}$ is a polynomial in $\mathbb{R}^{n}$ of degree $\left|\alpha\right|=\alpha_{1}+...+\alpha_{n}$ at most $\kappa-1$, and $x^{\alpha}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}...x_{n-1}^{\alpha_{n-1}}$. Let $d_{Q;\kappa}\equiv\dim L_{Q;\kappa}^{2}\left(\mu\right)$ be the dimension of the finite dimensional linear space $L_{Q;\kappa}^{2}\left(\mu\right)$. Moreover, for each $a\in\Gamma_{n}$, we may assume the wavelet $h_{Q;\kappa}^{a}$ is a translation and dilation of the unit wavelet $h_{Q_{0};\kappa}^{a}$, where $Q_{0}=\left[0,1\right)^{n}$ is the unit cube in $\mathbb{R}^{n}$.
Given a small positive constant $\eta>0$, define a smooth approximate identity by $\phi_{\eta}\left(x\right)\equiv\eta^{-n}\phi\left(\frac{x}{\eta}\right)$ where $\phi\in C_{c}^{\infty}\left(B_{\mathbb{R}^{n}}\left(0,1\right)\right)$ has unit integral, $\int_{\mathbb{R}^{n}}\phi\left(x\right)dx=1$, and vanishing moments of positive order less than $\kappa$, i.e.
(3.1) $\int\phi\left(x\right)x^{\gamma}dx=\delta_{\left|\gamma\right|}^{0}=\left\{\begin{array}[c]{ccc}1&\text{ if }&\left|\gamma\right|=0\\ 0&\text{ if }&0<\left|\gamma\right|<\kappa\end{array}\right.$.
The smooth Alpert ‘wavelets’ are defined by
$h_{Q;\kappa}^{a,\eta}\equiv h_{Q;\kappa}^{a}\ast\phi_{\eta\ell\left(Q\right)},$
and we have for $0\leq\left|\beta\right|<\kappa$,
$\int h_{Q;\kappa}^{a,\eta}\left(x\right)x^{\beta}dx=\int\phi_{\eta\ell\left(I\right)}\ast h_{Q;\kappa}^{a}\left(x\right)x^{\beta}dx=\int\int\phi_{\eta\ell\left(I\right)}\left(y\right)h_{Q;\kappa}^{a}\left(x-y\right)x^{\beta}dx$
$=\int\phi_{\eta\ell\left(I\right)}\left(y\right)\left\{\int h_{Q;\kappa}^{a}\left(x-y\right)x^{\beta}dx\right\}dy=\int\phi_{\eta\ell\left(I\right)}\left(y\right)\left\{\int h_{Q;\kappa}^{a}\left(x\right)\left(x+y\right)^{\beta}dx\right\}dy$
$=\int\phi_{\eta\ell\left(I\right)}\left(y\right)\left\{0\right\}dy=0,$
by translation invariance of Lebesgue measure.
There is a linear map $S_{\eta}^{\mathcal{D}}=S_{\kappa,\eta}^{\mathcal{D}}$, bounded and invertible on all $L^{p}\left(\mathbb{R}^{2}\right)$ spaces, $1<p<\infty$, such that if we define
$\bigtriangleup_{I;\kappa}^{\eta}f\equiv\left(\bigtriangleup_{I;\kappa}f\right)\ast\phi_{\eta\ell\left(I\right)},$
then
$\bigtriangleup_{I;\kappa}^{\eta}f\equiv\sum_{a\in\Gamma_{n}}\left\langle\left(S_{\eta}^{\mathcal{D}}\right)^{-1}f,h_{I;\kappa}^{a}\right\rangle h_{I;\kappa}^{a,\eta}=\sum_{a\in\Gamma_{n}}\left\langle\left(S_{\eta}^{\mathcal{D}}\right)^{-1}f,h_{I;\kappa}^{a}\right\rangle S_{\eta}^{\mathcal{D}}h_{I;\kappa}^{a}=\sum_{a\in\Gamma_{n}}\left(S_{\eta}^{\mathcal{D}}\bigtriangleup_{I;\kappa}\left(S_{\eta}^{\mathcal{D}}\right)^{-1}\right)f=\sum_{a\in\Gamma_{n}}\bigtriangleup_{I;\kappa}^{\spadesuit}f\ ,$
where $A^{\spadesuit}$ denotes the commutator $S_{\eta}^{\mathcal{D}}A\left(S_{\eta}^{\mathcal{D}}\right)^{-1}$ of an operator $A$ with $S_{\eta}^{\mathcal{D}}$.
Theorem 7 ([Saw7]).
Let $n\geq 2$ and $\kappa\in\mathbb{N}$ with $\kappa>\frac{n}{2}$. Then there is $\eta_{0}>0$ depending on $n$ and $\kappa$ such that for all $0<\eta<\eta_{0}$, and for all grids $\mathcal{D}$ in $\mathbb{R}^{n}$, and all $1<p<\infty$, there is a bounded invertible operator $S_{\eta}^{\mathcal{D}}=S_{\kappa,\eta}^{\mathcal{D}}$ on $L^{p}$, and a positive constant $C_{p,n,\eta}$ such that the collection of functions $\left\{h_{I;\kappa}^{a,\eta}\right\}_{I\in\mathcal{D},\ a\in\Gamma_{n}}$ is a $C_{p,n,\eta}$-frame for $L^{p}$, by which we mean,
(3.2) $f\left(x\right)$ $=\sum_{I\in\mathcal{D},\ a\in\Gamma_{n}}\bigtriangleup_{I;\kappa}^{\eta}f\left(x\right),\ \ \ \ \ \text{for all }f\in L^{p},$
$\text{where }\bigtriangleup_{I;\kappa}^{\eta}f$ $\equiv\sum_{a\in\Gamma_{n}}\left\langle\left(S_{\eta}^{\mathcal{D}}\right)^{-1}f,h_{I;\kappa}^{a}\right\rangle\ h_{I;\kappa}^{a,\eta}\ ,$
and with convergence of the sum in both the $L^{p}$ norm and almost everywhere, and
(3.3) $\frac{1}{C_{p,n,\eta}}\left\|f\right\|_{L^{p}}\leq\left\|\left(\sum_{I\in\mathcal{D}}\left|\bigtriangleup_{I;\kappa}^{\eta}f\right|^{2}\right)^{\frac{1}{2}}\right\|_{L^{p}}\leq C_{p,n,\eta}\left\|f\right\|_{L^{p}},\ \ \ \ \ \text{for all }f\in L^{p}.$
Moreover, the smooth Alpert wavelets $\left\{h_{I;\kappa}^{a,\eta}\right\}_{I\in\mathcal{D},\ a\in\Gamma_{n}}$ are translation and dilation invariant in the sense that $h_{I;\kappa}^{a,\eta}$ is a translate and dilate of the mother Alpert wavelet $h_{I_{0};\kappa}^{a,\eta}$ where $I_{0}$ is the unit cube in $\mathbb{R}^{n}$.
Notation 8.
We will often drop the index $a$ parameterized by the finite set $\Gamma_{n}$ as it plays no essential role in most of what follows, and it will be understood that when we write
$\bigtriangleup_{Q;\kappa}^{\eta}f=\left\langle\left(S_{\eta}^{\mathcal{D}}\right)^{-1}f,h_{Q;\kappa}\right\rangle h_{Q;\kappa}^{\eta}=\widehat{f}\left(Q\right)h_{Q;\kappa}^{\eta},$
we actually mean the Alpert pseudoprojection,
$\bigtriangleup_{Q;\kappa}^{\eta}f=\sum_{a\in\Gamma_{n}}\left\langle\left(S_{\eta}^{\mathcal{D}}\right)^{-1}f,h_{Q;\kappa}^{a}\right\rangle h_{Q;\kappa}^{\eta,a}=\sum_{a\in\Gamma_{n}}\widehat{f_{a}}\left(Q\right)h_{Q;\kappa}^{a,\eta}\ ,$
where $\widehat{f_{a}}\left(Q\right)$ is a convenient abbreviation for the inner product $\left\langle\left(S_{\eta}^{\mathcal{D}}\right)^{-1}f,h_{Q;\kappa}^{a}\right\rangle$ when $\kappa$ is understood. More precisely, one can view $\widehat{f}\left(Q\right)=\left\{\widehat{f_{a}}\left(Q\right)\right\}_{a\in\Gamma_{n}}$ and $h_{Q;\kappa}^{\eta}=\left\{h_{Q;\kappa}^{a,\eta}\right\}_{a\in\Gamma_{n}}$ as sequences of numbers and functions indexed by $\Gamma_{n}$, in which case $\widehat{f}\left(Q\right)h_{Q;\kappa}^{\eta}$ is the dot product of these two sequences. No confusion should arise between the Alpert coefficient $\widehat{g}\left(Q\right)$, $Q\in\mathcal{G}\left[U\right]$ and the Fourier transform $\widehat{g}\left(\xi\right)$, $\xi\in\mathbb{R}^{3}$, as the argument in the first is a square in $\mathcal{G}\left[U\right]$, while the argument in the second is a point in $\mathbb{R}^{3}$.
For $s\in\mathbb{N}$, and $K\in\mathcal{G}\left[S\right]$, with $S\subset\mathbb{R}^{2}$ centered at the origin as above, and with $\ell\left(K\right)\geq 2^{-s}$, we define the smooth Alpert pseudoprojection at scale $s$ by,
(3.4) $\mathsf{Q}_{s,K;\kappa}^{\eta}f\equiv\sum_{I\in\mathcal{G}_{s}\left[K\right]}\bigtriangleup_{I;\kappa}^{\eta}f\text{ and }f_{s,K}^{\Phi}\equiv\Phi_{\ast}\mathsf{Q}_{s,K;\kappa}^{\eta}f=\sum_{I\in\mathcal{G}_{s}\left[K\right]}\Phi_{\ast}\bigtriangleup_{I;\kappa}^{\eta}f.$
3.1. Initial setup and statement of the main Alpert characterization
Now we return to three dimensions. We recall some of the notation in [Saw7, Subsection 1.4] regarding local coordinates on the sphere, and pushforwards of smooth Alpert wavelets near the origin in $\mathbb{R}^{2}$. Fix a small cube $U_{0}$ in $\mathbb{R}^{n-1}$ with side length a negative power of $2$, and such that there is a translation $\mathcal{G}$ of the standard grid on $\mathbb{R}^{n-1}$ with the property that $U_{0}\in\mathcal{G}$, the grandparent $\pi_{\mathcal{G}}^{\left(2\right)}U_{0}$ of $U_{0}$ has the origin as a vertex, and $U_{0}$ is an interior grandchild of $U\equiv\pi_{\mathcal{G}}^{\left(2\right)}U_{0}$, so that
(3.5) $U_{0},U\in\mathcal{G}\text{ with }U_{0}\subset\frac{1}{2}U\text{.}$
Then parameterize a patch of the paraboloid $\mathbb{P}^{2}$ in the usual way, i.e. $\Phi:U\rightarrow\mathbb{S}^{2}$ by
$z=\Phi\left(x\right)\equiv\left(x,\left|x\right|^{2}\right)=\left(x_{1},x_{2},x_{1}^{2}+x_{2}^{2}\right).$
For $f\in L^{p}\left(U\right)$, define
$\mathcal{E}f\left(\xi\right)=\mathcal{E}_{U}f\left(\xi\right)\equiv\mathcal{F}\left(\Phi_{\ast}\left[f\left(x\right)dx\right]\right)=\int_{U}e^{-i\Phi\left(x\right)\cdot\xi}f\left(x\right)dx,$
where $\mathcal{F}$ is the Fourier transform in $\mathbb{R}^{3}$.
Recall that for $\nu>0$, we say that a triple $\left(U_{1},U_{2},U_{3}\right)$ of squares in $U\subset B_{\mathbb{R}^{2}}\left(0,\frac{1}{2}\right)$ is $\nu$-disjoint if
$\mathop{\rm diam}\left[\Phi\left(U_{k}\right)\right]\approx\mathop{\rm dist}\left[\Phi\left(U_{k}\right),\bigcup_{j:\ j\neq k}\Phi\left(U_{j}\right)\right]\geq\nu,\text{ for }1\leq k\leq 3,$
Definition 9.
Let $\varepsilon,\nu>0$ and $0<\delta<1$, $\kappa\in\mathbb{N}$ and $1<q<\infty$. Set $A\left(0,2^{r}\right)\equiv B\left(0,2^{r}\right)\setminus B\left(0,2^{r-1}\right)$ for $r\in\mathbb{N}$. Denote by $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ the trilinear smooth Alpert inequality,
(3.6) $\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}\leq C_{\varepsilon,\nu,\kappa,\delta,q}2^{\varepsilon r}\left\|f_{1}\right\|_{L^{\infty}}\left\|f_{2}\right\|_{L^{\infty}}\left\|f_{3}\right\|_{L^{\infty}}\ ,$
for all $r\in\mathbb{N}$, all $\nu$-disjoint triples $\left(U_{1},U_{2},U_{3}\right)$, all smooth Alpert pseudoprojections $\mathsf{Q}_{s_{k},U_{k}}^{\eta}$ with moment vanishing parameter $\kappa$ and with $s_{1}\leq s_{2}$ and $\frac{r}{1+\delta}<s_{2}\leq s_{3}<\frac{r}{1-\delta}$, and all $f_{1},f_{2},f_{3}\in L^{\infty}$.
Theorem 10.
Let $0<\delta<1$ and $\kappa>\frac{20}{\delta}$. The Fourier extension conjecture (1.1) for the paraboloid in $\mathbb{R}^{3}$ holds if and only if for every $q>3$ there is $\nu>0$ depending only on $q$, such that the disjoint smooth Alpert trilinear inequality $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ holds for all $\varepsilon>0$.
Note:
Integration on the left hand side of (3.6) is taken over the annulus $A\left(0,2^{s}\right)$ rather than the ball $B\left(0,2^{s}\right)$. Thus one should think of the $\mathcal{A}$ in $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa}\left(\otimes_{3}L^{p}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ as standing for ‘annulus’ as well as ‘Alpert’.
Note that the inequality (3.6) lies in the fully resonant spectrum, since
$\mathcal{E}\bigtriangleup_{I_{k};\kappa}^{\eta}f_{k}\left(\xi\right)=\sum_{I_{k}\in\mathcal{G}_{s}\left[U_{k}\right]}\widehat{f_{k}}\left(I_{k}\right)\int_{\mathbb{R}^{2}}e^{-i\Phi\left(x\right)\cdot\xi}h_{I_{k};\kappa}^{\eta}\left(x\right)dx\ ,$
where the wavelength of the oscillatory factor $e^{-i\Phi\left(x\right)\cdot\xi}$ on the support of $h_{I_{k};\kappa}^{\eta}$ is roughly $\frac{1}{\left|\xi\right|}\in\left(2^{-\left(1+\delta\right)s},2^{-\left(1-\delta\right)s}\right)$, and the side length of $I_{k}\,$is $2^{-s}$.
The smooth Alpert pseudoprojection $\mathsf{Q}_{s,U_{k}}^{\eta}=\sum_{I\in\mathcal{G}_{s}\left[U_{k}\right]}\bigtriangleup_{I;\kappa}^{\eta}$at level $s\in\mathbb{N}$ was introduced in [Saw7], where a probabilistic analogue of the Fourier extension conjecture was proved, but unlike the probabilistic analysis in [Saw7], which included a wide range of fully resonant inner products (from $2^{s}\,$to $2^{2s}$), the range of resonance in the disjoint smooth Alpert trilinear inequality $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta}\left(\otimes_{3}L^{q}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ is greatly reduced due to the convolution of disjoint patches on the sphere. It is crucial that we do not need to obtain estimates on how the constants depend on transversality in the trilinear result, something that is enabled by the argument of Bourgain and Guth [BoGu].
The disjoint smooth Alpert trilinear inequality $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta}\left(\otimes L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ represents the weakest formulation of the Fourier extension conjecture that the authors could find to date, requiring only smooth Alpert projections of bounded functions with the two largest scales near $s$ in a disjoint trilinear inequality, integration over the corresponding ‘small’ annulus $A\left(0,2^{s}\right)$ of resonance, and permitting the familiar small $\varepsilon$-power growth in $s$.
At the end of the paper we show how the probabilistic Fourier extension theorem in [Saw7] can be proved using a square function modification of the arguments in this paper, providing a new and arguably simpler proof.
3.2. Convolution of $\nu$-disjoint singular measures on the paraboloid
Let $\mu^{1}\equiv\Phi_{\ast}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}$ and $\mu^{2}\equiv\Phi_{\ast}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}$ denote singular measures on the paraboloid, that are pushforwards of smooth Alpert projections at levels $s_{1}<s_{2}$ of functions $f_{k}\in L^{p}\left(U_{k}\right)$, $1<p<\infty$, and where $\mathop{\rm diam}\left(U_{1}\right)\approx\mathop{\rm diam}\left(U_{2}\right)\approx\mathop{\rm dist}\left(U_{1},U_{2}\right)\gtrsim\nu>0$. For $z\in\mathbb{R}^{3}$, denote by $\omega_{z}$ the translate of a measure $\omega$ by $z$. We use duality to compute the convolution $\mu^{1}\ast\mu^{2}$ in terms of the measure-valued integral $\iiint_{w\in\mathbb{R}^{3}}\left[\mu_{w}^{1}\left(\cdot\right)\right]d\mu^{2}\left(w\right)$ as follows. For $F$ a continuous function on $\mathbb{R}^{3}$, write
$\left\langle F,\Phi_{\ast}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\ast\Phi_{\ast}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\right\rangle=\left\langle F,\mu^{1}\ast\mu^{2}\right\rangle=\left\langle F\left(\cdot\right),\iiint_{w\in\mathbb{R}^{3}}\mu_{w}^{1}\left(\cdot\right)d\mu^{2}\left(w\right)\right\rangle$
$=\iiint_{z\in\mathbb{R}^{3}}F\left(z\right)d\left[\iiint_{w\in\mathbb{R}^{3}}\mu_{w}^{1}\left(z\right)d\mu^{2}\left(w\right)\right]=\iiint_{w\in\mathbb{R}^{3}}\left\{\iiint_{z\in\mathbb{R}^{3}}F\left(z\right)d\mu_{w}^{1}\left(z\right)\right\}d\mu^{2}\left(w\right)$
$=\iiint_{w\in\mathbb{R}^{3}}\left\{\iiint_{z\in\mathbb{R}^{3}}F\left(z-w\right)d\mu^{1}\left(z\right)\right\}d\mu^{2}\left(w\right),$
and using the definitions of $\mu^{1}$ and $\mu^{2}$ as pushforwards respectively of $\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(v\right)dv$ and $\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(u\right)du$ by $\Phi$, we see that
$\left\langle F,\mu^{1}\ast\mu^{2}\right\rangle$ $=\iiint_{w\in\mathbb{R}^{3}}\left\{\iiint_{z\in\mathbb{R}^{3}}F\left(z-w\right)d\mu^{1}\left(z\right)\right\}d\mu^{2}\left(w\right)$
$=\iiint_{w\in\mathbb{R}^{3}}\left\{\iint_{v\in U_{1}}F\left(\Phi\left(v\right)-w\right)\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(v\right)dv\right\}d\mu^{2}\left(w\right)$
$=\iint_{u\in U_{2}}\iint_{v\in U_{1}}F\left(\Phi\left(v\right)-\Phi\left(u\right)\right)\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(v\right)dv\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(u\right)du.$
Taking limits we can let $F=\delta_{a}$, so that for $\left(v,u\right)\in U_{1}\times U_{2}$,
(3.7) $\left\langle\delta_{a},\mu^{1}\ast\mu^{2}\right\rangle$ $=\iiiint_{\left(v,u\right)\in U_{1}\times U_{2}}\delta_{a}\left(\Phi\left(v\right)-\Phi\left(u\right)\right)\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(v\right)dv\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(u\right)du$
$=\iiiint_{E_{a}}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(v\right)\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(u\right)\beta_{a}\left(u,v\right)dvdu,$
where $E_{a}$ is a line in $U_{1}\times U_{2}$ given by
$E_{a}$ $\equiv\left\{\left(v,u\right)\in U_{1}\times U_{2}:v-u=a^{\prime}\text{ and }\left|v\right|^{2}-\left|u\right|^{2}=a_{3}\right\}$
$=\left\{\left(u+a^{\prime},u\right):\left|u+a^{\prime}\right|^{2}-\left|u\right|^{2}=a_{3}\right\}=\left\{\left(u+a^{\prime},u\right):2a^{\prime}\cdot u=a_{3}-\left|a^{\prime}\right|^{2}\right\}$
$=\left\{\left(v,v-a^{\prime}\right):\left|v\right|^{2}-\left|v-a^{\prime}\right|^{2}=a_{3}\right\}=\left\{\left(v,v-a^{\prime}\right):2a^{\prime}\cdot v=a_{3}+\left|a^{\prime}\right|^{2}\right\},$
and $\beta_{a}\left(u,v\right)$ is the smooth density arising from the limiting passage from $F$ to $\delta_{a}$. This can be seen from an application of the implicit function theorem using $\left|a^{\prime}\right|=\left|v-u\right|\geq\nu>0$ by the $\nu$-disjoint assumption (or by direct calculation). We further conclude from the implicit function theorem that
[…]
where by adapted to scale $\delta>0$, we mean that $m^{th}$ order derivatives are bounded by $C_{m}\left(\frac{1}{\delta}\right)^{m}$. Moreover, the density $\mu^{1}\ast\mu^{2}\left(a\right)$’ of the absolutely continuous measure $\mu^{1}\ast\mu^{2}=\left(\mu^{1}\ast\mu^{2}\right)\left(a\right)da$ is given by
(3.9) $\mu^{1}\ast\mu^{2}\left(a\right)=\left\langle\delta_{a},\mu^{1}\ast\mu^{2}\right\rangle.$
At this point we note the crude inequality
$\left|\widehat{f}\left(I\right)\right|$ $=\left|\left\langle\left(S_{\kappa,\eta}^{\mathcal{G}}\right)^{-1}f,h_{I;\kappa}\right\rangle\right|\leq\left\|\left(S_{\kappa,\eta}^{\mathcal{G}}\right)^{-1}f\right\|_{L^{p}}\left\|h_{I;\kappa}\right\|_{L^{p^{\prime}}}$
$\lesssim\left\|\left(S_{\kappa,\eta}^{\mathcal{G}}\right)^{-1}\right\|_{L^{p}\rightarrow L^{p}}\left\|f\right\|_{L^{p}\left(U\right)}\left\|h_{I;\kappa}\right\|_{L^{p}}\left\|\mathbf{1}_{I}\right\|_{L^{p^{\prime}}}\lesssim\left\|f\right\|_{L^{p}\left(U\right)}\ell\left(I\right)^{-1}\ell\left(I\right)^{\frac{2}{p^{\prime}}}=\ell\left(I\right)^{\frac{1}{p^{\prime}}-\frac{1}{p}}\left\|f\right\|_{L^{p}}\ ,$
for any $1<p<\infty$, which gives
(3.10) $\left|\bigtriangleup_{I;\kappa}^{\eta}f\right|=\left|\widehat{f}\left(I\right)h_{I;\kappa}^{\eta}\right|\lesssim\ell\left(I\right)^{-\left(\frac{1}{p}-\frac{1}{p^{\prime}}\right)}\left\|f\right\|_{L^{p}}\ell\left(I\right)^{-1}\mathbf{1}_{\left(1+\eta\ell\left(I\right)\right)I}=\ell\left(I\right)^{-\frac{2}{p}}\left\|f\right\|_{L^{p}}\mathbf{1}_{\left(1+\eta\ell\left(I\right)\right)I},\ \ \ \ \ 1<p<\infty.$
Altogether (3.8), (3.9) and (3.10) prove the following lemma.
Lemma 11.
For $\mu^{1}$ and $\mu^{2}$ as above, i.e. $\mu^{1}\equiv\Phi_{\ast}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}$ and $\mu^{2}\equiv\Phi_{\ast}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}$ where $s_{1}<s_{2}$ and $f_{k}\in L^{p}\left(U_{k}\right)$, $1<p<\infty$, and $U_{1}$ and $U_{2}$ are $\nu$-separated. Then the following derivative estimates hold,
$\left|\nabla_{a}^{m}\left(\mu^{1}\ast\mu^{2}\right)\left(a\right)\right|$ $\lesssim C_{m}\frac{1}{\nu^{m}}2^{ms_{2}}2^{\frac{2}{p}\left(s_{1}+s_{2}\right)}\left\|f_{1}\right\|_{L^{p}\left(U_{1}\right)}\left\|f_{2}\right\|_{L^{p}\left(U_{2}\right)},$
$\text{for all }m$ $\geq 0\text{, }1<p<\infty\text{, and }a\in\mathop{\rm Supp}\left(\mu^{1}\ast\mu^{2}\right).$
Using Lemma 11 for $\nu$-disjoint singular measures on the paraboloid, we can now prove Theorem 10. Indeed, we will exploit the smoothness and moment vanishing properties of smooth Alpert wavelets to obtain geometric decay in convolutions of singular measures supported on the paraboloid.
Proof of Theorem 10.
Let $0<\delta<1$ and $\kappa>\frac{20}{\delta}$ as in the statement of Theorem 10. Without loss of generality we assume that $s_{1}<s_{2}<s_{3}$ and $R=2^{r}$, and we will repeatedly use the inequalities (3.8), (3.9) and (3.10) to prove Theorem 10 in three cases, followed by a wrapup. ∎
3.3. Case 1: $s_{3}$ is large
Proof continued.
First suppose that $s_{3}>\frac{r}{1-\delta}$. Then for $\left|\xi\right|\approx 2^{r}$, we have upon using $\kappa$-moment vanishing of $\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}$ and differentiating
$\exp_{\Phi,\xi}\left(x\right)\equiv e^{-i\Phi\left(x\right)\cdot\xi}=e^{-i\Phi\left(c_{I_{3}}\right)\cdot\xi}e^{-i\left[\Phi\left(x\right)-\Phi\left(c_{I_{3}}\right)\right]\cdot\xi}$
with respect to $x$, together with (3.10),
$\left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|=\left|\left[\Phi_{\ast}\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}\right]^{\wedge}\left(\xi\right)\right|=\left|\int e^{-iz\cdot\xi}\Phi_{\ast}\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}\left(z\right)dz\right|=\left|\int e^{-i\Phi\left(x\right)\cdot\xi}\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}\left(x\right)dx\right|$
$=\left|\int\left\{e^{-i\Phi\left(x\right)\cdot\xi}-\sum_{k=1}^{\kappa-1}\frac{\left[\left(x-c_{I_{3}}\right)\cdot\nabla\right]^{k}}{k!}\exp_{\Phi,\xi}\left(c_{I_{3}}\right)\right\}\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}\left(x\right)dx\right|$
$=\left|\int\left\{\frac{\left[\left(x-c_{I_{3}}\right)\cdot\nabla\right]^{\kappa}}{\kappa!}\exp_{\Phi,\xi}\left(\theta_{I_{3}}\right)\right\}\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}\left(x\right)dx\right|$
$\lesssim C_{\kappa}\int\left|2^{-s_{3}}\xi\right|^{\kappa}\left|\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}\left(x\right)\right|dx\lesssim C_{\kappa}\int\left(2^{-s_{3}}\left|\xi\right|\right)^{\kappa}2^{s_{3}\frac{2}{p}}\left\|f_{3}\right\|_{L^{p}}\mathbf{1}_{\left(1+\eta\ell\left(I_{3}\right)\right)I_{3}}dx$
$\leq C_{\kappa}2^{-\left(\kappa-\frac{2}{p}\right)s_{3}}2^{\kappa r}\left\|f_{3}\right\|_{L^{p}}=C_{\kappa}2^{-\kappa\left(s_{3}-r\right)}2^{\frac{2}{p}s_{3}}\left\|f_{3}\right\|_{L^{p}},$
and so
$\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim\left(\int_{A\left(0,2^{r}\right)}\left(\left\|f_{1}\right\|_{L^{1}}\left\|f_{2}\right\|_{L^{1}}C_{\kappa}2^{-\kappa\left(s_{3}-r\right)}2^{\frac{2}{p}s_{3}}\left\|f_{3}\right\|_{L^{p}}\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim C_{\kappa}2^{-\kappa\left(s_{3}-r\right)}2^{\frac{2}{p}s_{3}}\left(\int_{A\left(0,2^{r}\right)}d\xi\right)^{\frac{3}{q}}\left\|f_{1}\right\|_{L^{1}}\left\|f_{2}\right\|_{L^{1}}\left\|f_{3}\right\|_{L^{p}}$
$\lesssim C_{\kappa}2^{-\kappa\left(s_{3}-r\right)}2^{\frac{2}{p}s_{3}}2^{\frac{9}{q}r}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}$
$\leq C_{\kappa}2^{-\kappa\left(s_{3}-\left(1-\delta\right)s_{3}\right)}2^{\frac{2}{p}s_{3}}2^{\frac{9}{q}\left(1-\delta\right)s_{3}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}$
$=C_{\kappa}2^{-\kappa\delta s_{3}}2^{\left(\frac{9}{q}\left(1-\delta\right)+\frac{2}{p}\right)s_{3}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}\ .$
Since $\kappa>\frac{20}{\delta}$, $q>3$ and $p>1$, we have $\kappa\delta-\left(\frac{9}{q}\left(1-\delta\right)+\frac{2}{p}\right)\geq 15>1$. Then summing in $s_{3}>\frac{r}{1-\delta}$ and using $q>3$ in Minkowski’s inequality gives
(3.11) $\left(\int_{A\left(0,2^{r}\right)}\left(\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }s_{3}>\frac{r}{1-\delta}}\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }s_{3}>\frac{r}{1-\delta}}\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim C_{\kappa}\sum_{s_{3}>\frac{r}{1-\delta}}s_{3}^{2}2^{-s_{3}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}\lesssim C_{\kappa}r^{2}2^{-\frac{r}{1-\delta}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}.$
3.4. Case 2: $s_{2}$ is small
Proof continued.
Next we suppose $s_{1}<s_{2}\leq\frac{r}{1+\delta}$ and $s_{3}\leq\frac{r}{1-\delta}$, and $I_{1}\in\mathcal{G}_{s_{1}}\left[U_{1}\right]$ and $I_{2}\in\mathcal{G}_{s_{2}}\left[U_{2}\right]$. Then from Lemma 11, we obtain that
$F_{I_{1},I_{2}}\equiv\Phi_{\ast}\bigtriangleup_{I_{1};\kappa}^{\eta}f\ast\Phi_{\ast}\bigtriangleup_{I_{2};\kappa}^{\eta}f\left(z\right)$
is compactly supported in the three dimensional rectangle $2\left(\Phi\left(I_{1}\right)+\Phi\left(I_{2}\right)\right)$, and smoothly adapted to scale $\nu 2^{-s_{2}}$. Note that the smoothness scale from Lemma 11 is better than that obtained from the usual localization $\mathbf{1}_{A\left(0,2^{r}\right)}\leq\widehat{\varphi_{2^{-r}}}$, which is just $2^{-r}$ [^5] This smoothness $2^{-r}$ arising from localization is what was used in [Saw7], and the sharper smoothness $\nu 2^{-s_{2}}$ arising from $\nu$-disjointness used here is key to our alternate approach to the probabilistic Fourier extension theorem below.. Thus for $\xi\in A\left(0,2^{r}\right)$, we obtain from (3.10) that
$\left|\left[\Phi_{\ast}\bigtriangleup_{I_{1};\kappa}^{\eta}f\ast\Phi_{\ast}\bigtriangleup_{I_{2};\kappa}^{\eta}f\right]^{\wedge}\left(\xi\right)\right|=\left|\int e^{-iz\cdot\xi}F_{I_{1},I_{2}}\left(z\right)dz\right|\leq C_{N}\left(\frac{1}{\left|\xi\right|}\right)^{N}\int\left|\nabla^{N}F_{I_{1},I_{2}}\right|\left(z\right)dz$
$\leq C_{N}\left(\frac{2^{s_{2}}}{\nu\left|\xi\right|}\right)^{N}\left|\Phi\left(I_{1}\right)+\Phi\left(I_{2}\right)\right|\ell\left(I_{1}\right)^{-\frac{2}{p}}\left\|f_{1}\right\|_{L^{p}}\ell\left(I_{2}\right)^{-\frac{2}{p}}\left\|f_{2}\right\|_{L^{p}}$
$\approx C_{N}\nu\left(\frac{2^{s_{2}}}{\nu 2^{r}}\right)^{N}2^{-2s_{1}-s_{2}}2^{\frac{2}{p}s_{1}}2^{\frac{2}{p}s_{2}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\ ,$
since $\Phi\left(I_{1}\right)+\Phi\left(I_{2}\right)$ is roughly a three dimensional rectangle of dimensions $2^{-s_{1}}\times 2^{-s_{1}}\times\left(\sin\nu\right)2^{-s_{2}}$.
As a consequence we have
$\left(\int_{A\left(0,R\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}d\xi\right)^{\frac{3}{q}}$
$=\left(\int_{A\left(0,2^{r}\right)}\left(\left|\left[\Phi_{\ast}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f\ast\Phi_{\ast}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f\right]^{\wedge}\left(\xi\right)\right|\ \left|\widehat{\Phi_{\ast}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f}\left(\xi\right)\right|\right)^{\frac{q}{3}}d\xi\right)^{\frac{3}{q}}$
$\leq\sum_{I_{1}\in\mathcal{G}_{s_{1}}\left[U_{1}\right]}\sum_{I_{2}\in\mathcal{G}_{s_{2}}\left[U_{2}\right]}\left(\int_{A\left(0,2^{r}\right)}\left(\left|\left[\Phi_{\ast}\bigtriangleup_{I_{1};\kappa}^{\eta}f\ast\Phi_{\ast}\bigtriangleup_{I_{2};\kappa}^{\eta}f\right]^{\wedge}\left(\xi\right)\right|\ \left|\widehat{\Phi_{\ast}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f}\left(\xi\right)\right|\right)^{\frac{q}{3}}d\xi\right)^{\frac{3}{q}}$
$\leq\sum_{I_{1}\in\mathcal{G}_{s_{1}}\left[U_{1}\right]}\sum_{I_{2}\in\mathcal{G}_{s_{2}}\left[U_{2}\right]}\left(\int_{A\left(0,2^{r}\right)}\left(C_{N}\nu\left(\frac{2^{s_{2}}}{\nu 2^{r}}\right)^{N}2^{-2s_{1}-s_{2}}2^{\frac{2}{p}s_{1}}2^{\frac{2}{p}s_{2}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{1}}\right)^{\frac{q}{3}}d\xi\right)^{\frac{3}{q}},$
which is approximately,
$C_{N}\nu^{1-N}2^{2s_{1}}2^{2s_{2}}\left(\frac{2^{s_{2}}}{2^{r}}\right)^{N}2^{-2s_{1}-s_{2}}2^{\frac{2}{p}s_{1}}2^{\frac{2}{p}s_{2}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{1}}2^{r\frac{9}{q}}$
$\lesssim C_{N}\nu^{1-N}\left(\frac{2^{s_{2}}}{2^{r}}\right)^{N}2^{s_{2}}2^{\frac{2}{p}s_{1}}2^{\frac{2}{p}s_{2}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}2^{r\frac{9}{q}}$
$\leq C_{N}\nu^{1-N}2^{\left(\frac{9}{q}-N\right)r}2^{\left(N+1+\frac{4}{p}\right)s_{2}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}$
$<C_{N}\nu^{1-N}2^{\left(\frac{9}{q}-N\right)r}2^{\left(N+1+\frac{4}{p}\right)\frac{r}{1+\delta}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}\ .$
If we choose $\frac{N+1+\frac{4}{p}}{1+\delta}+\frac{9}{q}-N<-1$, then $2^{\left(\frac{9}{q}-N\right)r}2^{\left(N+1+\frac{4}{p}\right)\frac{r}{1+\delta}}<2^{-r}$ and so
$\sum_{s_{1},s_{2}=1}^{\frac{r}{1+\delta}}\left(\int_{A\left(0,R\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}d\xi\right)^{\frac{3}{q}}\lesssim C_{N}\nu^{1-N}r^{2}2^{-r}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}},$
for all $s_{3}\leq\frac{r}{1-\delta}$ and for $N$ sufficiently large depending only on $\delta$, $p$ and $q$, and in particular is independent of $s_{3}$.
Thus altogether we have
(3.12) $\left(\int_{A\left(0,2^{r}\right)}\left(\sum_{s_{1}\leq s_{2}\leq\frac{r}{1+\delta}\text{ and }s_{3}\leq\frac{r}{1-\delta}}\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\leq\sum_{s_{1}\leq s_{2}\leq\frac{r}{1+\delta}\text{ and }s_{3}\leq\frac{r}{1-\delta}}\left(\int_{B\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim\left(C_{N}\nu^{1-N}r^{2}2^{-r}\right)\frac{r}{1-\delta}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}\lesssim C_{N}\nu^{1-N}r^{3}2^{-r}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}.$
Remark 12.
Instead of appealing to Lemma 11 above, we could have fixed $\xi\in A\left(0,2^{r}\right)$, and then chosen a patch $U_{k}$ with $1\leq k\leq 2$ such that the unit vector $\xi^{\prime}\equiv\frac{\xi}{\left|\xi\right|}$ is at a positive angle depending on $\nu$, to the normal vectors of $\Phi\left(U_{k}\right)$, and then integrated by parts along the singular patch $\Phi\left(U_{k}\right)$. See e.g. [Saw7, Subsubsection 4.2.4] on tangential integration by parts for a similar calculation.
3.5. Wrapup of the proof
Proof continued.
Now we can finish the proof. With $f_{k}\in L^{\infty}\left(U_{k}\right)\subset L^{p}\left(U_{k}\right)$ and $R=2^{r}$,
$\left(\int_{A\left(0,R\right)}\left(\left|\mathcal{E}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$=\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\sum_{I_{1}\in\mathcal{G}\left[U_{1}\right]}\bigtriangleup_{I_{1};\kappa}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\sum_{I_{2}\in\mathcal{G}\left[U_{2}\right]}\bigtriangleup_{I_{2};\kappa}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\sum_{I_{3}\in\mathcal{G}\left[U_{3}\right]}\bigtriangleup_{I_{3};\kappa}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim\sum_{0\leq s_{1}\leq s_{2}\leq s_{3}}\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}},$
where the restriction to $I_{k}\in\mathcal{G}\left[U_{k}\right]$ can be made by adapting the reduction argument in the proof of [Saw7, Lemma 2 on page 5]. Next we obtain from (3.11) and (3.12), that the last line above is at most,
$\left\{\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }\frac{r}{1+\delta}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}}+\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }\frac{r}{1-\delta}\leq s_{3}}+\sum_{s_{1}\leq s_{2}\leq s_{3}\leq\frac{r}{1-\delta}\text{ and }s_{1}\leq s_{2}\leq\frac{r}{1+\delta}}\right\}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }\frac{r}{1+\delta}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}}\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +C_{\kappa,\nu}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}}.$
Recall that $0<\delta<1$ and $\kappa>\frac{20}{\delta}$ are given. Now by assumption, there is $\nu>0$ depending only on $q>3$ such that (3.6) holds with $2\delta$ and $\infty$ in place of $\delta$ and $p$ respectively, and this then gives
$\left(\int_{A\left(0,R\right)}\left(\left|\mathcal{E}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}\lesssim R^{\varepsilon}\left\|f_{1}\right\|_{L^{\infty}}\left\|f_{2}\right\|_{L^{\infty}}\left\|f_{3}\right\|_{L^{\infty}}.$
Finally, with $R=2^{r}$, we use that $B\left(0,R\right)$ is a union of the unit ball and at most $r$ annuli of the form $A\left(0,2^{t}\right)=B\left(0,2^{t}\right)\setminus B\left(0,2^{t-1}\right)$, in order to obtain,
$\int_{B\left(0,R\right)}\left(\left|\mathcal{E}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}d\xi$
$=\sum_{t=1}^{r}\int_{A\left(0,2^{t}\right)}\left(\left|\mathcal{E}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}d\xi+\int_{B\left(0,1\right)}\left(\left|\mathcal{E}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}d\xi$
$\lesssim\sum_{t=1}^{r}\left(2^{t\varepsilon}\left\|f_{1}\right\|_{L^{\infty}}\left\|f_{2}\right\|_{L^{\infty}}\left\|f_{3}\right\|_{L^{\infty}}\right)^{\frac{q}{3}}+C_{\nu}^{\frac{q}{3}}\left(\left\|f_{1}\right\|_{L^{\infty}}\left\|f_{2}\right\|_{L^{\infty}}\left\|f_{3}\right\|_{L^{\infty}}\right)^{\frac{q}{3}}$
$\lesssim C_{\nu}^{\frac{q}{3}}R^{\varepsilon\frac{q}{3}}\left(\left\|f_{1}\right\|_{L^{\infty}}\left\|f_{2}\right\|_{L^{\infty}}\left\|f_{3}\right\|_{L^{\infty}}\right)^{\frac{q}{3}},\ \ \ \ \ \text{for all }R\geq 1,$
which is $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$, where we have used $R^{\varepsilon}+C_{\nu}\leq C_{\nu}R^{\varepsilon}$ for $R^{\varepsilon},C_{\nu}\geq 1$. By Theorem 3, this is equivalent to the Fourier extension conjecture, and this completes the proof of Theorem 10. ∎
4. Appendix: Application to the probabilistic Fourier extension theorem
Here we briefly describe an alternate, and arguably simpler, approach to proving the three dimensional probabilistic Fourier extension inequality obtained in [Saw7]. However, this alternate approach has little likelihood of being extended to the Knapp segment at the boundary of allowable exponents, something that is not out of the question for the proof given in [Saw7].
The square function formulation of the probabilistic inequality proved in [Saw7] is,
(4.1) $\left\|\mathcal{S}_{\mathop{\rm Fourier}}f\right\|_{L^{q}\left(\lambda_{n}\right)}\lesssim\left\|f\right\|_{L^{q}\left(B\left(0,\frac{1}{2}\right)\right)}\ ,$
where $\mathcal{S}_{\mathop{\rm Fourier}}$ is the Fourier square function defined by
(4.2) $\mathcal{S}_{\mathop{\rm Fourier}}f\equiv\left(\sum_{I\in\mathcal{G}\left[U\right]}\left|\mathcal{E}\mathbf{1}_{U_{0}}\bigtriangleup_{I;\kappa}^{n-1,\eta}f\right|^{2}\right)^{\frac{1}{2}}.$
Here the extension operator $\mathcal{E}$ is defined by
(4.3) $\mathcal{E}f\left(\xi\right)\equiv\int_{B_{2}\left(0,\frac{1}{2}\right)}e^{-i\Phi\left(x\right)\cdot\xi}f\left(x\right)dx,\ \ \ \ \ \xi\in\mathbb{R}^{n},$
for $f\in L^{p}\left(B_{2}\left(0,\frac{1}{2}\right)\right)$. Thus $\mathcal{E}f=\mathcal{F}\Phi_{\ast}\left(f\lambda_{2}\right)=\widehat{\Phi_{\ast}\left(f\lambda_{2}\right)}$, where $\Phi_{\ast}\nu$ denotes the pushforward of a measure $\nu$ under the map $\Phi$, and $\lambda_{2}$ is Lebesgue measure in the plane.
Here we conjecture the square function analogue of the second part of Theorem 10.
Definition 13.
Suppose $n=3$. Let $\varepsilon,\nu>0$, $0<\delta<1$, and $1<q<\infty$. Denote by $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta,\mathop{\rm square}}\left(\otimes_{3}L^{q}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ the disjoint smooth Alpert square function trilinear inequality
$\left\|\mathcal{S}_{\mathop{\rm Fourier}}\mathsf{Q}_{U_{1}}^{s_{1}}f_{1}\ \mathcal{S}_{\mathop{\rm Fourier}}\mathsf{Q}_{U_{2}}^{s_{2}}f_{2}\ \mathcal{S}_{\mathop{\rm Fourier}}\mathsf{Q}_{U_{3}}^{s_{3}}f_{3}\right\|_{L^{\frac{q}{3}}\left(A\left(0,2^{s}\right)\right)}\leq C_{\varepsilon,\nu}2^{\varepsilon s}\left\|f_{1}\right\|_{L^{\infty}\left(U\right)}\left\|f_{2}\right\|_{L^{\infty}\left(U\right)}\left\|f_{3}\right\|_{L^{\infty}\left(U\right)}\ ,$
taken over all $s_{1}\leq s_{2}$ and $s_{2},s_{3}\in\left(\left(1-\delta\right)s,\left(1+\delta\right)s\right)$, all $f_{k}\in L^{q}\left(U_{k}\right)$, and all triples $\left(U_{1},U_{2},U_{3}\right)\subset U^{3}$ that satisfy the weak $\nu$-disjoint condition,
$\mathop{\rm diam}\left[\Phi\left(U_{k}\right)\right]\approx\mathop{\rm dist}\left[\Phi\left(U_{k}\right),\bigcup_{j:\ j\neq k}\Phi\left(U_{j}\right)\right]\gtrsim\nu,\text{ for }1\leq k\leq 3,$
Conjecture 14.
Let $0<\delta<1$ and $\kappa>\frac{10}{\delta}$. The square function inequality (4.1) holds in $\mathbb{R}^{3}$ if and only if for every $q>3$ there is $\nu>0$ such that the disjoint smooth Alpert square function trilinear inequality $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta,\mathop{\rm square}}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ holds for all $\varepsilon>0$.
It should be possible to prove this square function variant by tracing through the proof of Theorem 10 above and making modifications for the square function. See for example [RiSa2], where these modifications are carried out for ‘father’ wavelets $\varphi_{I}$ smoothly adapted to a square II, and their projections $\bigtriangleup_{I}f\equiv\left\langle f,\varphi_{I}\right\rangle\varphi_{I}$. However, an additional difficulty arises for smooth Alpert wavelets $h_{I;\kappa}^{\eta}$, since the pseudoprojections $\bigtriangleup_{I;\kappa}^{\eta}f\equiv\left\langle S_{\kappa,\eta}^{-1}f,h_{I;\kappa}\varphi_{I}\right\rangle h_{I;\kappa}^{\eta}$ involve an operator $S_{\kappa,\eta}^{-1}$ that is bounded only on $L^{p}$ for $1<p<\infty$, and not on $L^{\infty}$. In this case, parabolic rescalings must be performed with $f\in L^{p}\left(U\right)$, which introduces an additional growth factor $C_{p}2^{\frac{2}{p}s}$, that turns out to be harmless because one can take $p$ arbitrarily large. As mentioned earlier, the proof of this conjecture will be addressed in a future paper.
Conjecture 14 can be used to give an alternate proof of the probabilistic analogue of Fourier extension in [Saw7]. In fact, the case $n=3$ of Proposition 34 in [Saw7] says that for $q>3$, there is $\varepsilon_{q}>0$ such that for every $s\in\mathbb{N}$, and every $f\in L^{q}\left(U\right)$, we have,
(4.4) $\mathbb{E}_{\pm}\left\|\mathcal{E}\left[\left(\pm\mathsf{Q}_{U}^{s}\right)^{\spadesuit}f\right]\right\|_{L^{q}\left(B\left(0,2^{s}\right)\right)}\lesssim 2^{-s\varepsilon_{q}}\left\|f\right\|_{L^{q}\left(U\mathbb{R}^{n-1}\right)},$
where $\mathbb{E}_{\pm}$ denotes the average over the ‘martingale transforms’ $\left(\pm\mathsf{Q}_{U}^{s}\right)^{\spadesuit}f$, and the implied constant depends on $q$ and $U$, but is independent of $s\in\mathbb{N}$. Indeed, (4.4) is easily obtained by computing the norms when $p=2$ and $4$, and then using the expectation $\mathbb{E}_{\pm}$ to eliminate a large number of off diagonal terms in the $L^{4}$ estimate, resulting in a geometric decay in $s$ that exactly balances the growth in $s$ for the $L^{2}$ estimate when $q=3$. Moreover, we can enlarge the ball $B\left(0,2^{s}\right)$ to $B\left(0,2^{\left(1+\delta\right)s}\right)$ provided $\delta\leq C\varepsilon$. See [Saw7, Section 5] for details.
Then using Khintchine’s inequality, we obtain the square function formulation of (4.4), namely that for $q>3$, there is $\varepsilon_{q}>0$ such that for every $s\in\mathbb{N}$, and every $f\in L^{q}\left(U\right)$, we have,
$\left\|\mathcal{S}_{\mathop{\rm Fourier}}\mathsf{Q}_{U}^{s}f\right\|_{L^{q}\left(B\left(0,2^{\left(1+\delta\right)s}\right)\right)}\lesssim 2^{-s\varepsilon_{q}}\left\|f\right\|_{L^{p}\left(U\right)},\ \ \ \ \ s\in\mathbb{N}.$
Thus we conclude that $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta,\mathop{\rm square}}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ holds, even with a negative exponent $\varepsilon$. Now Conjecture 14 completes the alternate proof of the square function formulation (4.1) of the probabilistic Fourier extension theorem in the case $n=3$.